### v1

# AutoRev: Automatic Peer Review System for Academic Research Papers

Abstract
Generating a review for an academic research paper is a complex task that requires a deep understanding of the document’s content and the interdependencies between its sections. It demands not only insight into technical details but also an appreciation of the paper’s overall coherence and structure. Recent methods have predominantly focused on fine-tuning large language models (LLMs) to address this challenge. However, they often overlook the computational and performance limitations imposed by long input token lengths. To address this, we introduce AutoRev, an Automatic Peer Review System for Academic Research Papers. Our novel framework represents an academic document as a graph, enabling the extraction of the most critical passages that contribute significantly to the review. This graph-based approach demonstrates effectiveness for review generation and is potentially adaptable to various downstream tasks, such as question answering, summarization, and document representation. When applied to review generation, our method outperforms SOTA baselines by an average of 58.72% across all evaluation metrics. We hope that our work will stimulate further research in applying graph-based extraction techniques to other downstream tasks in NLP. We plan to make our code public upon acceptance.
AutoRev: Automatic Peer Review System for Academic Research Papers
Maitreya Prafulla Chitale1thanks: Equal contribution.  Ketaki Mangesh Shetye [^1] footnotemark: 1 Harshit Gupta1Manav Chaudhary1Vasudeva Varma1 1IIIT Hyderabad {maitreya.chitale, ketaki.shetye, harshit.g, manav.chaudhary}@research.iiit.ac.in vv@iiit.ac.in
1 Introduction
Figure 1: Example graph constructed out of a paper consisting of 2 Headings, 2 Sub-Headings, 4 Passages, and 6 Sentences.
Recent research has increasingly focused on developing automatic systems for peer review, including tools that assist reviewers, assess peer-review processes, and directly generate reviews for academic papers Yu et al. (2024); Aitymbetov and Zorbas (2025); Church et al. (2025). Concurrently, there has been growing interest in examining the strengths and limitations of such systems, particularly the use of large language models (LLMs) for automated review generation Ye et al. (2024). Motivated by both of these trends, we propose AutoRev, a system designed to generate peer reviews from a guidance perspective, offering constructive feedback to support researchers and reviewers.
With the rise of Graph Neural Networks, the use of graphs as extractors or as guides for facilitating downstream tasks has recently gained traction in the NLP community Chitale et al. (2025); Gao et al. (2025); Agarwal et al. (2025). Inspired by this, we propose to use graphs in the domain of academic paper analysis, specifically, generating reviews that closely replicate those produced by existing peer review systems in conferences. AutoRev can generate high-quality reviews without relying on any prior human evaluations of the paper. Furthermore, by leveraging the graph’s capability as an extractor of passages, we significantly reduce the input token length to the LLM. Reducing the token length of the input has been shown to enhance LLM performance Hsieh et al. (2024).
To summarize our main contributions:
  • We propose a system for generating concise, structured reviews (Summary, Strengths, Weaknesses, and Questions) to assist authors and reviewers.
  • To our knowledge, this is the first approach using graphs in this domain.
  • Our approach improves over baselines by 58.72% on average across all metrics, without relying on human-written reviews and staying efficient.
It is important to highlight that this paper does not seek to determine the final reviews of a submission. Rather, we present our framework as a tool to provide timely feedback to researchers and serve as a valuable guide for reviewers.
2 Related Works
A significant amount of research has been conducted on aiding peer reviews for academic papers. Some studies explore the involvement of AI in the selection process Aitymbetov and Zorbas (2025), while some studies work on generating meta-reviews Kumar et al. (2023). Sufficient discussion and avenues have been explored where AI/NLP techniques can contribute towards aiding automatic peer reviews Kuznetsov et al. (2024). Though, to the best of our knowledge, significant research has not been done in the field of generating a whole review for an academic paper.
A recent study by Faizullah et al. (2024) explores a DPR-based approach for generating suggestive limitations. It highlights the necessity of using DPR to address length constraints and extract only the most relevant passages from a paper. This study served as an initial motivation for incorporating DPR into our approach.
Another notable work by Yu et al. (2024) introduces an automated paper reviewing framework. The work also includes a standardized dataset of GPT-4 OpenAI et al. (2024) consolidated reviews collected from various academic venues. To enhance review generation, the authors propose a self-correction strategy during LLM fine-tuning to ensure a higher quality of the generated reviews. We adopt this work as our baseline and aim to develop a more computationally efficient approach than direct LLM fine-tuning while achieving higher-quality reviews. The dataset publicised by the authors of this study is being directly used by us in our work.
Ye et al. (2024) highlight key issues in LLM-generated reviews, such as giving overly high ratings in some cases and being easily influenced by authors to generate only minor limitations. Considering the contributions of this work, we reinforce the need for our approach, not as a tool for final review decisions but as a means to provide constructive feedback to researchers and reviewers.
3 Dataset
The dataset used in our study is borrowed from Yu et al. (2024), comprising papers submitted to various conferences along with multiple peer reviews per paper. Among these, only the datasets for ICLR 2024 and NeurIPS 2023 are publicly available. However, the number of GPT-4-consolidated reviews for NeurIPS 2023 is limited to just 235 samples. Hence, our analysis focuses only on the ICLR 2024 papers. As reported in Yu et al. (2024), the dataset statistics for the ICLR 2024 papers are summarized in Table 1.
Attribute Value
Total Papers 5,653
Average Tokens per Paper 9,815
Total Reviews 21,839
Acceptance Rate 37%
Domain Machine Learning
Table 1: Dataset statistics for the ICLR 2024 dataset, as reported in Yu et al. (2024).
Each paper has multiple peer reviews including sections such as Summary, Strengths, Weaknesses, Questions, Soundness, Presentation, Contribution, Confidence, and Rating to assess the quality and clarity of the paper.
From the dataset, we utilize only the Summary, Strengths, Weaknesses, and Questions sections from each review, along with the full content of the corresponding paper. Other sections, such as Ratings are excluded, as our framework is designed to provide constructive feedback rather than produce a final decision. The selected sections offer qualitative assessments of the paper’s content and provide valuable insights to guide both authors and reviewers. Additionally, we use the GPT-4 OpenAI et al. (2024) consolidated reviews dataset, sourced from Yu et al. (2024), for finetuning LLMs to generate final structured reviews. This dataset covers 20% of randomly selected ICLR-24 papers (2036 papers), where the multiple peer reviews are consolidated using GPT-4 to generate integrated reviews in a unified format.
4 Methodology
In this section, we describe the implementation of our proposed framework. Section 4.1 describes the initial passage extraction using Dense Passage Retrieval (DPR) Karpukhin et al. (2020). These extracted passages are used as training data in Section 4.3, to train a Graph Neural Network (GNN) that learns to extract relevant passages from a research paper. This trained GNN enables our model to perform competitively on the test set without the need for grounding on human-generated reviews. Post that, we describe our process of fine-tuning LLMs to generate the final reviews, in Section 4.5. An overview of our methodology can be found in Figure 2
Figure 2: The system begins with a parsed document from which DPR extracts passages for GNN training. The GNN learns to identify key passages relevant to review generation. The trained GNN selects important passages without requiring human-generated reviews. These passages, combined with a structured prompt, are input to fine-tune an LLM. The fine-tuned LLM generates the final review efficiently, without external grounding.
4.1 DPR-based Passage Extraction
Dense Passage Retrieval (DPR) is a neural retrieval model, primarily trained on question-answering datasets, that encodes both queries and context into dense vector representations. Given an input query, DPR retrieves the most relevant passages based on their semantic similarity in the dense embedding space. We use a pre-trained DPR model, to identify the most relevant passages corresponding to the different sections of the reviews that are Summary, Strengths, Weaknesses, and Questions, for each paper in the training set. Specifically, we use the context encoder [^1] https://huggingface.co/facebook/dpr-ctx_encoder-multiset-base and the query encoder [^2] https://huggingface.co/facebook/dpr-question_encoder-multiset-base from the multiset-base DPR variant.
[…]
This final set of passages represents the most relevant excerpts from the full paper that are semantically linked to the review content. These extracted passages serve as the gold-standard outputs for training the GNN, as described in Section 4.3
4.2 Graph Construction for Representing a Paper
The idea of graph construction employed in our work is inspired by Zhao et al. (2024). We borrow their idea of representing a document as a hierarchical graph but extend it beyond the sentence level. We propose to build a deeper graph, which contributes towards learning more intricate inter-dependencies of passages and sentences in a paper.
Specifically, we formalize the document structure as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ represents nodes and $\mathcal{E}$ denotes edges encoding hierarchical and sequential relationships.
Node Definitions: Our graph construction is designed to be generalizable across diverse academic paper structures, accommodating arbitrary levels of nesting and heterogeneous section organization.
  • Paper Node ($v_{\text{paper}}$): A single root node representing the entire document.
  • Heading Nodes $\mathcal{H}={h_{1},h_{2},\dots,h_{m}}$: Nodes corresponding to top-level section headings.
  • Sub-heading Nodes $\mathcal{S}_{i}={s_{i1},s_{i2},\dots,s_{ik}}$: Nodes representing sub-headings under each heading $h_{i}$. This structure naturally extends to accommodate multiple levels of nesting, making it suitable for documents with deeper hierarchical organization.
[…]
  • Sentence Nodes $\mathcal{T}_{jk}={t_{jk1},t_{jk2},\dots,t_{jkp}}$: Nodes for individual sentences within each passage $o_{jk}$, tokenized using NLTK Bird and Loper (2004).
Edge Definitions: Edges capture both hierarchical relationships and sequential flow within the document.
  • Root connections:
    $(v_{\text{paper}},h_{i})\in\mathcal{E}_{\text{hier}},\ \forall h_{i}\in\mathcal{H}$
  • Heading-subheading:
    $(h_{i},s{ij})\in\mathcal{E}_{\text{hier}},\ \forall s{ij}\in\mathcal{S}i$
  • Content links:
    $(s_{ij},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ if }\mathcal{S}_{i}\neq\emptyset;\\ (h_{i},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ otherwise}$
  • Passage-sentence:
    $(o_{jk},t_{jkq})\in\mathcal{E}_{\text{hier}},\ \forall t_{jkq}\in\mathcal{T}_{jk}$
  • Sequential edges (horizontal flow):
    • Passages: $(o_{jk},o_{j(k+1)})\in\mathcal{E}_{\text{seq}},\ \forall\ 1\leq k<n$
    • Sentences: $(t_{jkq},t_{jk(q+1)})\in\mathcal{E}_{\text{seq}},\ \forall\ 1\leq q<p$
The final edge set is the union of hierarchical and sequential edges:
$\mathcal{E}=\mathcal{E}_{\text{hier}}\cup\mathcal{E}_{\text{seq}}$ (1)
where $\mathcal{E}_{\text{hier}}$ captures the vertical (structural) relationships and $\mathcal{E}_{\text{seq}}$ captures the horizontal (textual progression) flow. This graph construction is agnostic to section depth and flexibly adapts to any nested structure within academic documents. A toy example of such a graph, built from a paper with 2 Headings, 2 Sub-Headings, 4 Passages, and 6 Sentences, is shown in Figure 1.
4.3 GNN Training for Passage Retrieval
To train a GNN on the graph constructed in Section 4.2, we take the data generated in Section 4.1, which maps each paper $p_{i}$ to a set of important passages, $\mathcal{O}_{i}=\{o_{1},o_{2},\dots,o_{n}\}$. Each passage is important for generating a review for the paper it is associated with. This mapping is used as a training set for training the GNN.
The node embeddings for sentence nodes are initialized with embeddings from a Sentence-Transformer Reimers and Gurevych (2019) model, namely all-mpnet-base-v2. The node embeddings for all other nodes are initialized with zero embeddings of dimension $768$, matching that of the Sentence-Transformer model.
Now, a Graph Attention Network Veličković et al. (2018) is used to train the graph to predict the important passages. After passing the initial representation of the graph nodes through three layers of GATConv [^3] https://pytorch-geometric.readthedocs.io/en/latest/generated/torch_geometric.nn.conv.GATConv.html (PyTorch Geometric’s implementation of Graph Attention Networks), we apply a non-linear classifier to the final passage node representations. The classification is defined as follows:
  • A label of 0 indicates that the passage is not important for review generation.
  • A label of 1 indicates that the passage is important for review generation.
The GNN is trained using various configurations based on the values of $k$ and $m$, as described in the DPR process in Section 4.1. These trained GNN models are denoted as GNN $(k,m)$.
We hypothesize that our choice of graph construction allows our model to capture the intricate inter-dependencies of sentences and passages across different sections within an academic paper, and effectively make a choice about which passage(s) is/are important. The dependencies, according to our hypothesis, will be learned from the sequential and hierarchical connections that we have introduced in our graph construction method. The passing of information from one node to another will be enhanced by the GATConv (Equations 2, 3, and 4) layers, which involves the concept of attention to selectively propagate messages in the Message Passing layer of a GNN.
$\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$ (2)
$e_{ij}=\text{LeakyReLU}\Bigl{(}\mathbf{a}^{T}\bigl{[}\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}\bigr{]}\Bigr{)}$ (3)
$\alpha_{ij}=\frac{\exp\bigl{(}e_{ij}\bigr{)}}{\sum_{k\in\mathcal{N}(i)}\exp\bigl{(}e_{ik}\bigr{)}}$ (4)
Additional implementation details related to this section are available in Appendix B.
Figure 3: Excerpt of a test sample where passages highlighted in cyan colour refer to relevant passages identified by GNN only and yellow colour refer to relevant passages identified by both GNN and DPR
As illustrated in an example from our test set (Figure 3), the DPR-based approach fails to retrieve a key passage outlining the main contributions in the paper, whereas GNN successfully identifies it. Thus, leveraging sequentially and hierarchically connected graph nodes and GAT in GNN proves to be more effective in identifying important passages that help to enhance the review generation process. The GNN extracted passages, along with a structured prompt, serve as the input to LLM fine-tuning, as outlined in Section 4.5.
4.4 Grounding Requirements and Inference Efficiency
The advantage of training a GNN for this important classification task is twofold. First, unlike DPR-based approaches which require grounding at inference time via access to human-generated reviews or synthesized queries, our GNN-based method operates in a fully autonomous manner. Once trained, the GNN identifies salient passages directly from the graph-structured representation of the paper, eliminating the need for any external review or human guidance during inference. This distinction is crucial for real-world deployment, where human reviews may not be available or feasible to use.
Second, by selecting only the passages deemed most relevant, we significantly reduce the input length and associated computational costs for downstream LLM-based review generation. This not only leads to higher performance of LLMs, as outlined by Hsieh et al. (2024) but also leads to higher scores of the generated reviews (compared to our baselines), as reflected in our performance metrics (see Section 5). The statistics related to the decrease in number of passages and token length post-extraction by the GNN Model are reported in Table 2.
Input to LLM Avg Token Len Avg Passages
Full Paper 9445.92 47.07
GNN $(5,5)$ 4147.15 29.49
GNN $(5,3)$ 2899.59 21.00
Table 2: Average token length (Avg. Token Len) and Average passage count (Avg. Passages) for a full paper, our best-configuration: GNN $(5,5)$, and our fastest-configuration: GNN $(5,3)$.
4.5 Fine-Tuning Large Language Models for Review Generation
General-purpose LLMs are not trained specifically to generate peer reviews for academic research papers. Our motivation for fine-tuning is to ensure that the models learn to generate qualitatively sound, factual and structured peer reviews.
We select Mistral-Instruct-v0.2 (Mistral) Jiang et al. (2023), inspired by our baseline SEA-EA, which also uses Mistral-Instruct-v0.2 as its underlying backbone model. We also use Llama 3.1 8B (Llama) Dubey et al. (2024), as it is an open-weight state-of-the-art (SOTA) model with strong text generation capabilities for a variety of tasks. We employ LoRA (Low-Rank Adaptation) Hu et al. (2022), for fine-tuning due to its efficiency in adapting large language models to particular tasks while maintaining computational feasibility. This allows us to leverage the pre-trained capabilities of Llama and Mistral while efficiently adapting them for review generation.
We use the trained GNN, as described in Section 4.3, to retrieve relevant passages for each paper that serve as the LLM’s input during fine-tuning. The GPT-4 consolidated reviews for the ICLR 2024 papers, as released by Yu et al. (2024), are regarded as gold standard references for training the LLM to generate final reviews. The dataset split is summarized in Table 3.
Dataset Split Number of Samples
Training Set 1,765
Test Set 271
Table 3: Number of samples used for training and testing sourced from Yu et al. (2024)
Both Mistral and Llama are fine-tuned on the passages retrieved, using various configurations of the trained GNN $(k,m)$, as described in Section 4.3. These fine-tuned models are denoted Mistral $(k,m)$ and Llama $(k,m)$, respectively. The prompt and hyperparameter settings used during fine-tuning are provided in Appendix C.
5 Results and Analysis
Model ROUGE F1 ROUGE Recall BERTScore-F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
SEA-E 34.29 11.71 14.98 23.28 7.90 10.13 84.04
SEA-EA 36.97 13.02 15.88 25.62 8.97 10.97 84.15
Mistral $(5,3)$ 53.44 15.47 22.39 53.18 15.42 22.31 86.44 4.34
Llama $(5,3)$ 54.24 15.91 22.59 53.75 15.77 22.41 86.59 3.60
Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
Table 4: Overall performance on the ICLR-24 reviews test dataset, comparing baselines from Yu et al. (2024) with our fine-tuned models. For both Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral), the configuration with $(5,5)$ achieves the best performance, while $(5,3)$ has the shortest training time. Bold highlights the best scores, and underlined values indicate the second-best. Please note that the GPU configurations for finetuning Llama and Mistral were different. For details, refer to Appendix C.
We evaluate the quality of the generated reviews using ROUGE-1, ROUGE-2, and ROUGE-L (both Recall and F1) Lin (2004), as well as BERTScore-F1 Zhang* et al. (2020), on the test set of GPT-4-consolidated reviews for ICLR 2024 papers. ROUGE measures similarity by analyzing n-grams, with recall indicating how well the generated reviews capture key information from the original papers and the F1 score balancing precision and recall. In contrast, BERTScore assesses semantic similarity in the embedding space, providing a deeper understanding of meaning beyond surface-level text matching.
5.1 Performance Comparison with Baselines
We compare the baselines SEA-E and SEA-EA, two automated review generation approaches, as introduced by Yu et al. (2024). As mentioned in Table 4, our results show that we consistently outperform the baselines in all metrics, with particularly significant improvements in the ROUGE metrics (Recall and F1). Our best-performing configuration, Llama $(5,5)$ (selected based on the highest BERTScore), achieves an improvement of 58.72% over the state-of-the-art (SOTA) baselines across all evaluation metrics. To ensure that these gains are not solely due to using a more capable LLM, we also evaluated our method using Mistral. The results demonstrate that even with Mistral, integrating our GNN-based passage selection leads to substantial improvements, confirming that the performance boost is primarily due to our graph-based retrieval strategy rather than the choice of LLM alone.
5.2 Performance of Our Model
Model Review Section ROUGE-F1 ROUGE-Recall BERTScore-F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
Mistral $(5,3)$ Questions 39.24 9.33 20.00 40.90 9.74 20.91 86.15 4.34
Weaknesses 40.12 9.28 19.56 38.83 8.98 18.98 86.75
Strengths 43.15 11.17 22.07 43.63 11.25 22.47 87.83
Summary 45.18 13.08 25.28 45.35 13.14 25.38 88.09
Llama $(5,3)$ Questions 40.15 9.66 20.02 39.53 9.47 19.85 86.26 3.60
Weaknesses 40.74 9.56 19.63 40.59 9.50 19.60 86.79
Strengths 42.93 11.13 22.09 43.36 11.20 22.44 87.84
Summary 45.65 13.60 25.85 46.67 13.92 26.45 88.23
Mistral $(5,5)$ Questions 39.20 9.00 19.91 40.67 9.32 20.83 86.09 6.44
Weaknesses 40.27 9.21 19.46 39.49 9.01 19.13 86.71
Strengths 43.45 11.49 22.34 44.11 11.61 22.82 87.84
Summary 46.24 14.44 26.32 46.56 14.57 26.48 88.33
Llama $(5,5)$ Questions 39.54 9.31 20.01 38.43 9.02 19.62 86.25 5.30
Weaknesses 40.72 9.66 19.79 40.72 9.59 19.93 86.77
Strengths 43.43 11.76 22.24 43.36 11.69 22.34 87.91
Summary 46.92 14.90 26.74 47.30 14.99 26.92 88.36
Table 5: The sectionwise performance across the ICLR-24 reviews test dataset after finetuning Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral). Per section of the review, the best results have been made bold, while the second-best results have been underlined. Please note that the GPU configurations for finetuning Llama and Mistral were different. For details, refer to Appendix C.
As shown in Table 5, our best-performing configuration, Llama $(5,5)$ consistently outperforms the baselines across ROUGE-F1, ROUGE-Recall, and BERTScore metrics. Llama $(5,5)$ configuration is finetuned with input passages inferred from the GNN $(5,5)$. This highlights its ability to generate reviews that closely align with gold-standard references. The integration of a Graph Attention Network (GAT) enables the model to effectively capture complex dependencies between relevant passages and sentences within the paper, offering the essential context required to generate comprehensive reviews.
Notably, even the fastest variant of our fine-tuned models, Llama $(5,3)$, which requires the shortest training time, performs competitively. Despite using a smaller number of retrieved passages, this variant shows only an average marginal decrease of 0.71% in performance metrics compared to the best-performing configuration, Llama $(5,5)$. Further, it consistently outperforms the state-of-the-art (SOTA) baselines, while reducing training time by 32.07%. This trend is consistently observed across all metrics for Mistral too. It highlights that our GNN-based approach can achieve near-optimal performance with significantly less computational overhead.
As shown in Table 5, a detailed analysis of our model’s performance across different review sections reveals that the Summary and Strengths sections achieve higher scores across ROUGE and BERTScore in our best-performing model, Llama $(5,5)$. This indicates that the model is effective at capturing and utilizing essential context from the paper to generate these key components of the review. Conversely, the Weaknesses and Questions sections receive comparatively lower scores. This indicates that identifying limitations and raising questions require a deeper semantic understanding of the paper as weaknesses may not always be explicitly stated in the passages of the paper. Overall, GNN helps to reflect on the content of the paper comprehensively to generate suitable reviews.
5.3 LLM-as-a-judge Evaluation
Review Review Quality Assessment Criteria
Reviewer Confidence Review Thoroughness Review Constructiveness Review Helpfulness
Gold 3.98 3.87 3.79 3.79
Mistral $(5,3)$ 3.50 3.35 3.44 3.32
Llama $(5,3)$ 3.61 3.45 3.47 3.39
Mistral $(5,5)$ 3.48 3.36 3.43 3.34
Llama $(5,5)$ 3.63 3.44 3.47 3.41
Table 6: The table reports the qualitative evaluation of model-generated reviews using Gemini-2.0-Flash as an LLM judge. It presents the average scores (out of 5) across four review quality criteria: Reviewer Confidence, Review Thoroughness, Review Constructiveness, and Review Helpfulness, in comparison to gold reviews on the test set. Bold highlights the model best scores, and underlined values indicate the second-best.
As mentioned by Zheng et al. (2023), the LLM-as-a-judge method offers an explainable and affordable way to approximate human preferences, which are otherwise costly to gather. We employ a large language model, Gemini-2.0-Flash [^4] https://deepmind.google/technologies/gemini/ as an automated judge to evaluate the quality of reviews generated by our proposed model as compared to the gold-standard reviews. The evaluation criteria are designed to capture key aspects of review quality and include:
  1. 1.
    Reviewer Confidence: reflects the reviewer’s depth of technical understanding and engagement with the content.
  2. 2.
    Review Thoroughness: measures how well the review covers the methodology, results, and reproducibility.
  3. 3.
    Review Constructiveness: assesses the clarity and usefulness of suggestions for improvement.
  4. 4.
    Review Helpfulness: indicates how well the review identifies important issues and explains them clearly for the authors.
The model is prompted to assign a score from 1 to 5 based on the above criteria. The evaluation prompt is detailed in Appendix D. As shown in Table 6, our best-performing model, Llama $(5,5)$, scores only 9.59% below gold reviews across all criteria. The fastest variant, Llama $(5,3)$, achieves comparable quality with just a 9.78% drop relative to gold reviews while reducing training time by 32.07%. All Llama and Mistral variants consistently score above 3.0, indicating that the reviews generated grasp key contributions and provide relevant feedback with actionable suggestions.
5.4 Ablation Studies
We perform ablation studies by systematically varying the values of $k\in\{1,3,5\}$ and $m\in\{3,5\}$ during LLM fine-tuning. The detailed results are reported in Appendix A. The results indicate that increasing $k$ and $m$ provides richer context to both Mistral and Llama to generate reviews, leading to consistent improvements across all evaluation metrics. Additionally, we observe that higher values of $m$ result in longer training times, as the GNN retrieves and passes a larger set of passages as input for finetuning the LLMs.
6 Conclusion
In this work, we introduce AutoRev, an automatic peer review generation system designed specifically for academic research papers. Our approach leverages a novel graph-based framework to represent the structure of a paper which enables the effective extraction of important passages using a Graph Neural Network. This strategy of selecting passages and reducing the input length to LLMs addresses the limitations of traditional methods, like finetuning using the entire paper as input Yu et al. (2024), which leads to sub-optimal performance of LLMs due to large input token length Hsieh et al. (2024).
Unlike DPR-based methods that require initial grounding on human written references to retrieve relevant passages, AutoRev employs a GNN that learns to identify important content directly from the paper’s graphical representation, without the need for grounding on existing reviews. This not only streamlines the process but also reduces computational overhead while enhancing review quality.
Our experiments on the ICLR 2024 dataset demonstrate that AutoRev consistently outperforms existing baselines, by 58.72% on popular evaluation metrics like ROUGE, and BERTScore.
Overall, AutoRev provides a promising solution for pre-review feedback, aiding both authors and reviewers in the academic publication process. Future work will explore extending this framework to a broader range of academic domains and investigating further improvements in graph construction and neural retrieval techniques to continue advancing the state-of-the-art in automated review generation.
Limitations
While our framework provides a robust and high-performing solution for review generation, we acknowledge that no automatic system can fully replicate human judgment in complex tasks such as peer review, as noted by Ye et al. (2024). Our goal is to assist authors in refining their papers and we aim for our system to support human reviewers. Currently, our system is primarily focused on ICLR papers due to the limited availability of public datasets. However, we believe in the generalization capabilities of our model, given the proven effectiveness of graph-based approaches in various downstream NLP tasks, as discussed in Section 1.
Ethical Considerations
The deployment of AutoRev raises key ethical concerns related to bias, misuse, and academic integrity. Since the system is trained on existing peer reviews, it may inherit and amplify biases present in those texts, such as leniency toward well-known authors or under-representation of certain research areas, potentially skewing evaluations. Moreover, there is a risk that users may over-rely on AutoRev for making or substituting final review decisions, which could undermine the depth and critical thinking central to human peer review. Additionally, the realistic quality of AI-generated reviews opens the possibility of unethical use, such as submitting AI-authored critiques as original human work. To address these issues, we advocate for responsible usage of AutoRev as a supplementary tool, accompanied by transparency in AI involvement and safeguards to preserve academic integrity.
Appendix A Ablation studies
The quantitative performance of different combinations of $k$ (number of review sentences concatenated; $k=1,3,5$) and $m$ (number of top DPR-retrieved passages used to train the GNN; $m=3,5$) for review generation is reported in Table 7 and Table 8, evaluated against the gold-standard reviews. While calculating the ROUGE metrics, we enable the use_stemmer option to ensure consistent matching of words with different morphological forms, to improve the reliability of the evaluation.
Model ROUGE-F1 ROUGE-Recall BERTScore F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
Mistral $(1,3)$ 53.54 15.81 22.68 53.26 15.74 22.59 86.51 7.60
Mistral $(1,5)$ 53.54 15.76 22.64 52.81 15.56 22.36 86.51 10.22
Mistral $(3,3)$ 53.86 15.80 22.62 53.21 15.63 22.36 86.54 6.33
Mistral $(3,5)$ 53.87 15.91 22.74 53.43 15.79 22.57 86.54 7.64
Mistral $(5,3)$ 53.44 15.47 22.39 53.18 15.42 22.31 86.44 4.34
Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
Table 7: Overall performance of Mistral across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
Model ROUGE-F1 ROUGE-Recall BERTScore F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
Llama $(1,3)$ 54.42 16.19 22.76 53.71 15.98 22.48 86.61 6.65
Llama $(1,5)$ 54.40 16.09 22.82 53.58 15.86 22.50 86.61 8.23
Llama $(3,3)$ 54.52 16.03 22.67 53.79 15.83 22.39 86.61 5.11
Llama $(3,5)$ 54.61 16.30 22.88 53.78 16.06 22.56 86.66 6.28
Llama $(5,3)$ 54.24 15.91 22.59 53.75 15.77 22.41 86.59 3.60
Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
Table 8: Overall performance of Llama across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
Appendix B GNN Implementation Details
We trained and tested our model on a single NVIDIA RTX6000 GPU with 48GB of VRAM.
B.1 Model Architecture
The model contains a Graph Neural Network (GNN) module and a passage classifier.
GNN Module
The GNN uses three graph attention (GATConv) layers. The first GATConv layer processes input features (of size 768 (embedding dimension of Sentence Encoder)) into 128-dimensional features with 4 attention heads, producing a concatenated output of $128\times 4$ dimensions. A residual connection is incorporated by applying a linear layer to project the input to $128\times 4$, which is added to the GATConv output. This is followed by an ELU activation and dropout ($p=0.2$).
The second GATConv layer maps the $128\times 4$ input to 256-dimensional features using 2 attention heads (output dimension $256\times 2$). A residual connection similarly projects the prior layer’s output to $256\times 2$, followed by ELU and dropout ($p=0.2$). The final GATConv layer upscales the $256\times 2$ features to the target output dimension (1024 in our case) without residual connections and dropout layers.
Passage Classifier
For passage nodes, we employ a feedforward classifier to map the output embeddings from the GNN to the final label space. The classifier is implemented as a multi-layer perceptron (MLP) comprising the following sequence of operations:
  • A linear transformation projecting the input to a 512-dimensional space.
  • A SiLU activation followed by dropout with a probability of $p=0.3$.
  • A second linear layer projecting to 256 dimensions.
  • Layer normalization applied to the 256-dimensional features.
  • Another SiLU activation.
  • A final linear layer projecting to the number of target classes (typically 2 for binary classification).
This architecture enables the model to learn nonlinear transformations of the graph-encoded passage representations while incorporating regularization and normalization for stable training.
B.2 Training
The model is trained for 30 epochs, with the best validation performance checkpoint retained for inference. The loss function used for training is the Cross Entropy Loss.
Appendix C Finetuning Implementation Details
We fine-tune Llama-3.1-8B using two NVIDIA L40S GPUs and Mistral using a single NVIDIA RTX 6000 Ada GPU. Both GPUs are equipped with 48GB of GDDR6 VRAM.
We apply Low-Rank Adaptation (LoRA) with a rank of 8 and a scaling factor ($\alpha=16$). LoRA is integrated into the following transformer layers: q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj. Models are fine-tuned for 3 epochs. The total training time for different configurations is reported in Table 7 and Table 8. The full set of hyperparameters used during training is listed in Table 9.
Hyperparameter Value
Batch Size 1
Gradient Accumulation Steps 1
16-bit Floating Point Precision True
Optimizer adamw_torch
Learning Rate $1\times 10^{-4}$
Max Gradient Norm 0.3
Warmup Ratio 0.03
LoRA Rank 8
LoRA scaling factor 16
LoRA Dropout 0.05
Weight Decay 0.01
LR Scheduler Type cosine
Table 9: Hyperparameters Used During Finetuning
Prompt for Finetuning
The prompt used for finetuning both Llama 3.1 8B and Mistral-Instruct-v0.2 can be found below:
Prompt Used During Finetuning Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. ### Instruction:
Generate a structured review for the research
paper passages provided below. The review should include a summary of the paper, its strengths, weaknesses, and questions for the authors. Consider that the paper is being reviewed for the ICLR conference. ### Research Paper Passages:
### Review for the paper:
[…]
The extracted relevant passages are inserted ahead of the "Research Paper Passages" heading. Similarly, the corresponding content for each review section (Summary, Strengths, Weaknesses, and Questions) is added after their respective headings in the generated review.
Appendix D LLM-as-a-judge evaluation
** Instruction ** : You are an expert meta-reviewer. Evaluate the following peer review text according to the scoring
[…]
Give rating for the categories described below on the full scale from 1(lowest rating) to 5(highest rating).
Assign scores independently per category (they may differ significantly).
**Category 1: Reviewer Confidence**
1 – Unqualified: Admits lack of expertise | Fundamental misunderstandings | Fails to engage with technical content
2 – Limited: Surface-level comprehension | Misses >2 key concepts | Confuses established methods
3 – Functional: Grasps main contributions | Identifies 1-2 technical flaws | Limited literature context
4 – Qualified: Precise technical critique | Contextualizes in recent work (>3 citations) |
Identifies methodological tradeoffs
5 – Authoritative: Demonstrates field leadership | Predicts follow-up research directions |
Reveals fundamental limitations in approach
**Category 2: Review Thoroughness**
1 – Negligent: <3 substantive comments | Skips methodology/results | Fails to address claims
2 – Patchy: Only addresses obvious aspects | 0 analysis of experiments | No ethical consideration
3 – Standard: Checks main claims | Basic methodology review | 1 reproducibility concern noted
4 – Rigorous: Line-by-line methodology check | Statistical validity analysis | 2+ improvement suggestions
5 – Forensic: Alternative approach proposals | Re-analysis of key results |
Compliance checklist (ethics, reproducibility, etc.)
**Category 3: Constructiveness**
1 – Destructive: Personal attacks | Unsubstantiated dismissal | 0 actionable feedback
2 – Demanding: Unrealistic requests ("redo all experiments") | Vague quality complaints
3 – Directive: General improvement areas | 1-2 specific examples | Mixed positive/negative tone
4 – Strategic: Priority-ranked suggestions | Alternative methodology paths | Conference-specific improvements
5 – Transformative: Step-by-step revision roadmap | Code/experiment snippets provided | Before-after examples
**Category 4: Helpfulness**
1 – Misleading: Factually incorrect suggestions | Contradicts paper’s goals | Harmful recommendations
2 – Abstract: "Improve writing" without examples | "Add experiments" without design
3 – Guided: Section-specific edits | 2-3 implementable suggestions | Partial error diagnosis
4 – Impactful: Directly address weaknesses | Templates/examples provided | Clear priority levels
5 – Revelatory: Exposes fundamental flaws | Provides benchmark comparisons | Supplies missing literature
**Scoring Methodology**
For EACH category:
1. Find the HIGHEST level (5-1) where ALL criteria are FULLY satisfied (Base Level)
2. Check criteria in the NEXT HIGHER LEVEL (Base Level + 1):
   - Calculate partial credit: (Satisfied Criteria in Higher Level / Total Criteria in Higher Level)
3. Final Score = Base Level + Partial Credit
4. Round to 2 decimal places (0.33 increments for 3-criteria levels, 0.5 for 2-criteria levels)
[…]
Respond ONLY with this JSON format:
  "review_scores": {{
    "Category 1": "",
    "Category 2": "",
    "Category 3": "",
    "Category 4": ""
Figure 4: Prompt used during evaluating the quality of generated reviews with LLM as a Judge
The prompt used for LLM-as-a-judge with Gemini-2.0-Flash is shown in Figure 4. The corresponding model-generated review is inserted under the appropriate subheadings within the prompt.
To demonstrate the effectiveness and sensitivity of our evaluation prompt, we include two manually crafted sample reviews: one representing a poorly written, low-quality review, and another simulating an average review. These examples are not generated by our system but are intentionally designed to test whether the LLM-as-a-judge can distinguish between different qualities of feedback. This validation step supports the reliability of the evaluation framework in scoring model outputs.
Example 1: Structured Critical Review (Moderate Quality)
Summary: The paper tackles a topic that lacks clear novelty, presents a poorly motivated approach, and delivers incomplete experiments with unconvincing results. The writing is disorganized, and key methodological and comparative details are missing, making it difficult to assess the contribution.
Strengths:
  • Addresses a relevant area of graph-based modeling (though similar to existing work).
  • Attempts to combine embeddings and graph structure in a single framework.
Weaknesses:
  • Motivation is vague and fails to establish why the problem matters.
  • No thorough comparison to closely related methods; literature review is superficial.
  • Methodology section is under-specified, lacking algorithmic details and hyperparameter settings.
  • Experimental results are incomplete: missing error bars, statistical analysis, and ablation studies.
  • Figures and tables are low resolution or poorly labeled, hindering interpretability.
Questions:
  • What is the core novelty compared to the most recent graph-embedding methods?
  • How were key hyperparameters chosen, and how sensitive are the results to those choices?
  • Can you provide ablation studies showing the contribution of each component in the model?
  • Why are error bars and statistical significance tests omitted from the reported results?
  • How does your method scale to larger datasets, and what is its computational complexity?
The evaluation scores obtained for Example-1 are as follows,
  1. 1.
    Reviewer Confidence : 3.0
  2. 2.
    Review Thoroughness : 3.0
  3. 3.
    Review Constructiveness : 3.33
  4. 4.
    Review Helpfulness : 3.0
Example 2: Informal and Unstructured Review (Low Quality)
This paper is pretty weak. The writing is sloppy and hard to follow. I’m not convinced the idea is novel. There are lots of papers out there doing similar stuff. The experiments seem incomplete, and the results aren’t very impressive. I don’t think this will be of interest to ICLR readers. You should do more experiments or maybe scrap the whole thing and start over.
Comments:
  • The motivation is unclear.
  • There’s no thorough comparison to related work.
  • The methodology description is too brief.
  • Results tables have no error bars.
  • Figures are low resolution and hard to read.
Recommendation: Reject.
The evaluation scores obtained for Example-2 are as follows,
  1. 1.
    Reviewer Confidence : 2.0
  2. 2.
    Review Thoroughness: 1.0
  3. 3.
    Review Constructiveness : 2.0
  4. 4.
    Review Helpfulness : 2.0