### v2

# AutoRev: Multi-Modal Graph Retrieval for Automated Peer-Review Generation

Abstract.
Enhancing the quality and efficiency of academic publishing is critical for both authors and reviewers, as research papers are central to scholarly communication and a major source of high-quality content on the web. To support this goal, we propose AutoRev, an automatic peer-review system designed to provide actionable, high-quality feedback to both reviewers and authors. AutoRev leverages a novel Multi-Modal Retrieval-Augmented Generation (RAG) framework that combines textual and graphical representations of academic papers. By modelling documents as graphs, AutoRev effectively retrieves the most pertinent information, significantly reducing the input context length for LLMs and thereby enhancing their review generation capabilities. Experimental results show that AutoRev outperforms state-of-the-art baselines by up to 58.72% and demonstrates competitive performance in human evaluations against ground truth reviews. We envision AutoRev as a powerful tool to streamline the peer-review workflow, alleviating challenges and enabling scalable, high-quality scholarly publishing. By guiding both authors and reviewers, AutoRev has the potential to accelerate the dissemination of quality research on the web at a larger scale. Code will be released upon acceptance.
Graph Neural Networks (GNNs), Retrieval Augmented Generation (RAG), Large Language Models (LLMs), Automatic Peer Reviewing
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Information extractionccs: Computing methodologies Natural language generationccs: Information systems Language modelsccs: Information systems Document structure
1. Introduction
Figure 1. Example document graph with nodes for headings, sub-headings, passages, and sentences. This structure allows AutoRev to capture hierarchical and sequential dependencies across sections.:
Recent research has increasingly focused on developing automatic systems for peer review, including tools that assist reviewers, assess peer-review processes, and directly generate reviews for academic papers (Yu et al., 2024; Aitymbetov and Zorbas, 2025; D’Arcy et al., 2024; Idahl and Ahmadi, 2025). Concurrently, there has been growing interest in examining the strengths and limitations of such systems, particularly the use of large language models (LLMs) for automated review generation (Ye et al., 2024). Motivated by both of these trends, we propose AutoRev, a system designed to generate peer reviews from a guidance perspective, offering constructive feedback to support researchers and reviewers.
With the rise of Graph Neural Networks, the use of graphs as extractors or as guides for facilitating downstream tasks has recently gained traction in the NLP community (Chitale et al., 2025; Gao et al., 2025; Agarwal et al., 2025; Zhu et al., 2025). Inspired by this, we propose to use graphs in the domain of academic paper analysis, specifically, in generating reviews that closely replicate the structure and quality of those produced by existing peer review systems in conferences. Furthermore, by leveraging the graph’s capability as an extractor of passages, we significantly reduce the input token length to the LLM. Reducing the token length of the input is shown to enhance LLM performance (Hsieh et al., 2024). To summarize our main contributions:
  1. (1)
    We propose a system for generating concise, structured reviews (Summary, Strengths, Weaknesses, and Questions) to assist authors and reviewers. To our knowledge, this is the first approach which uses graphs in this domain.
  2. (2)
    Our approach improves over SoTA baselines by 58.72% on average across all metrics while being computationally efficient.
  3. (3)
    We conduct thorough human evaluations and other automatic evaluations to demonstrate the competitive performance of our method against ground truth reviews.
It is important to highlight that this paper does not seek to determine the final reviews of a paper. Rather, we present our framework as a tool to provide timely feedback to researchers and serve as a valuable guide for reviewers.
We envision AutoRev as a step towards scalable, automated peer-review systems that can complement human judgment. With the continued surge in conference submissions across disciplines, the demand for such tools is increasingly pressing [^1] https://aaai.org/aaai-launches-ai-powered-peer-review-assessment-system/. By alleviating the bottlenecks of the traditional peer-review process, AutoRev can help accelerate the dissemination of high-quality research on the web, ensuring that credible scholarly knowledge remains accessible at scale.
2. Related Works
A significant amount of research has been conducted on aiding peer reviews for academic papers. Some studies explore the involvement of AI in the reviewer selection process (Aitymbetov and Zorbas, 2025), while others focus on generating meta-reviews (Kumar et al., 2023). Prior work has also examined how AI/NLP techniques can contribute to various aspects of the peer-review pipeline (Kuznetsov et al., 2024).
Faizullah et al. (2024) explore a DPR-based approach for generating suggestive limitations, highlighting the necessity of retrieval methods to overcome length constraints. Building on this, we use DPR outputs as the training data for our Graph Neural Network (GNN), enabling it to learn to identify salient passages for review generation. Yu et al. (2024) introduce an automated paper reviewing framework (SEA) that standardizes reviews using GPT-4222https://arxiv.org/abs/2303.08774 consolidation and proposes a self-correction strategy during fine-tuning. Ye et al. (2024) further caution that LLM-generated reviews may overrate papers or produce superficial feedback, reinforcing our design goal of providing a structured and constructive review rather than final accept/reject decisions.
Among the most relevant recent systems, MARG (D’Arcy et al., 2024) employs a multi-agent framework, distributing different sections of a paper to specialized GPT-4 agents. This enables the model to handle long papers and produce more specific comments, but it comes at the cost of heavy computation and frequent inter-agent coordination errors. OpenReviewer (Idahl and Ahmadi, 2025) instead fully fine-tunes Llama-3.1-8B-Instruct on 79k reviews from ICLR and NeurIPS to generate structured reviews, adhering to the conference guidelines. This method of large-scale full fine-tuning is very costly.
We adopt SEA (Yu et al., 2024) as our primary baseline since it is closest in scope to our work, representing a recent state-of-the-art framework for automated peer review generation on the same ICLR dataset. Other systems such as MARG and OpenReviewer differ substantially in design. MARG uses multi-agent GPT-4 orchestration, while OpenReviewer relies on large-scale fine-tuning, making direct benchmarking less meaningful. Thus, we benchmark against SEA for a fair and reproducible comparison.
The integration of graph-based techniques into Retrieval Augmented Generation (RAG) (Lewis et al., 2020) is an emerging research area that aims to enhance the reasoning and factual grounding of LLMs. A recent comprehensive survey by Zhu et al. (2025) provides a foundational taxonomy for this field, categorizing systems based on their database construction, retrieval algorithms, and processing pipelines. Within this framework, AutoRev is a domain-specific Graph-RAG application that generates a hierarchical graph directly from a scientific paper’s text. It employs a learning-based Graph Attention Network (GAT) (Veličković et al., 2018) for retrieval that reduces the context length of input to LLMs, reducing the computational costs.
3. Dataset
The dataset used in our study is borrowed from Yu et al. (2024), comprising papers submitted to various conferences along with multiple peer reviews per paper. Among these, the ICLR 2024 and NeurIPS 2023 datasets are publicly released, though the GPT-4–consolidated reviews for NeurIPS 2023 are not released. Hence, our analysis focuses only on the ICLR 2024 papers. As reported in Yu et al. (2024), the dataset statistics for the ICLR 2024 papers are summarized in Table 1.
Table 1. Dataset statistics for the ICLR 2024 dataset, as reported in Yu et al. (2024).
Attribute Value
Total Papers 5,653
Average Tokens per Paper 9,815
Total Reviews 21,839
Acceptance Rate 37%
Domain Machine Learning
Each paper has multiple peer reviews including sections such as Summary, Strengths, Weaknesses, Questions, Soundness, Presentation, Contribution, Confidence, and Rating to assess the quality and clarity of the paper.
From the dataset, we utilize only the Summary, Strengths, Weaknesses, and Questions sections from each review, along with the full content of the corresponding paper. Other sections, such as Ratings are excluded, as our framework is designed to provide constructive feedback rather than produce a final decision. The selected sections offer qualitative assessments of the paper’s content and provide valuable insights to guide both authors and reviewers. For our finetuning process, we utilize the GPT-4 consolidated reviews dataset introduced by Yu et al. (2024). This dataset is constructed by using GPT-4 to consolidate multiple human-written peer reviews from OpenReview [^3] https://openreview.net/ into a single, unified review for each paper.
4. Methodology
Figure 2. The system begins with a parsed document from which DPR extracts passages for GNN training. The GNN learns to identify key passages relevant to review generation. The trained GNN selects important passages without requiring human-generated reviews. These passages, combined with a structured prompt, are input to fine-tune an LLM. The fine-tuned LLM generates the final review efficiently, without external grounding.:
In this section, we describe the implementation of our proposed framework. Section 4.1 describes the initial passage extraction using Dense Passage Retrieval (DPR) (Karpukhin et al., 2020). The extracted passages are used as training data in Section 4.3, to train a Graph Neural Network (GNN) that learns to extract relevant passages from a research paper. The trained GNN enables our model to perform competitively on the test set without the need for grounding on human-generated reviews, as discussed in Section 4.4. Post that, we describe our process of fine-tuning LLMs to generate the final reviews, in Section 4.5. An overview of our methodology can be found in Figure 2.
4.1. DPR-based Passage Extraction
Dense Passage Retrieval (DPR) is a neural retrieval model, primarily trained on question-answering datasets, that encodes both queries and context into dense vector representations. Given an input query, DPR retrieves the most relevant passages based on their semantic similarity in the dense embedding space. We use a pre-trained DPR model, to identify the most relevant passages corresponding to the different sections of the reviews that are Summary, Strengths, Weaknesses, and Questions, for each paper in the training set. Specifically, we use the context encoder [^4] https://huggingface.co/facebook/dpr-ctx_encoder-multiset-base and the query encoder [^5] https://huggingface.co/facebook/dpr-question_encoder-multiset-base from the multiset-base DPR variant.
[…]
This final set of passages represents the most relevant excerpts from the full paper that are semantically linked to the review content. These extracted passages serve as the ground truth labels for training the GNN, as described in Section 4.3
4.2. Graph Construction for Representing a Paper
The idea of graph construction employed in our work is inspired by Zhao et al. (2024). We borrow their idea of representing a document as a hierarchical graph but extend it beyond sentence level. We propose to build a deeper graph, which contributes towards learning more intricate inter-dependencies of passages and sentences in a paper.
Specifically, we formalize the document structure as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where nodes $\mathcal{V}$ represent text segments with semantically encoded features, and $\mathcal{E}$ denote the hierarchical and sequential relationships that form the document’s structure. Our graph construction is designed to be generalizable across diverse academic paper structures, accommodating arbitrary levels of nesting and heterogeneous section organization.
4.2.1. Node Definitions
  • Paper Node ($v_{\text{paper}}$): A single root node representing the entire document.
  • Heading Nodes ($\mathcal{H}=\{h_{1},h_{2},\dots,h_{m}\}$): Nodes corresponding to top-level section headings.
  • Sub-heading Nodes ($\mathcal{S}_{i}=\{s_{i1},s_{i2},\dots,s_{ik}\}$): Nodes representing sub-headings under each heading $h_{i}$. This structure naturally extends to accommodate multiple levels of nesting, making it suitable for documents with deeper hierarchical organization.
[…]
  • Sentence Nodes ($\mathcal{T}_{jk}=\{t_{jk1},t_{jk2},\dots,t_{jkp}\}$): Nodes for individual sentences within each passage $o_{jk}$, tokenized using NLTK (Bird and Loper, 2004).
4.2.2. Edge Definitions
Edges capture both hierarchical and sequential relationships within the document.
  • Root Connections: $(v_{\text{paper}},h_{i})\in\mathcal{E}_{\text{hier}}$ for all $h_{i}\in\mathcal{H}$.
  • Heading-SubHeading Connections:
    $(h_{i},s_{ij})\in\mathcal{E}_{\text{hier}}$ for all $s_{ij}\in\mathcal{S}_{i}$.
  • Content Links:
    $(s_{ij},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ if }\mathcal{S}_{i}\neq\emptyset;(h_{i},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ otherwise.}$
  • Passage-Sentence Connections:
    $(o_{jk},t_{jkq})\in\mathcal{E}_{\text{hier}}$ for all $t_{jkq}\in\mathcal{T}_{jk}$.
  • Sequential Edges:
    For passages: $(o_{jk},o_{j(k+1)})\in\mathcal{E}_{\text{seq}}$ for $1\leq k<n$.
    For sentences: $(t_{jkq},t_{jk(q+1)})\in\mathcal{E}_{\text{seq}}$ for $1\leq q<p$.
The final edge set is the union of hierarchical and sequential edges:
(1) $\mathcal{E}=\mathcal{E}_{\text{hier}}\cup\mathcal{E}_{\text{seq}},$
where $\mathcal{E}_{\text{hier}}$ captures the vertical (structural) relationships and $\mathcal{E}_{\text{seq}}$ captures the horizontal (textual progression) flow. This graph construction is agnostic to section depth and flexibly adapts to any nested structure within academic documents. A toy example of such a graph, built from a paper with 2 Headings, 2 Sub-Headings, 4 Passages, and 6 Sentences, is shown in Figure 1.
4.3. GNN Training for Multi Modal Passage Retrieval
Figure 3. Excerpt of a test sample where passages highlighted in cyan colour refer to relevant passages identified by GNN only and yellow colour refer to relevant passages identified by both GNN and DPR:
To train a GNN on the graph constructed in Section 4.2, we take the data generated in Section 4.1, which maps each paper $p_{i}$ to a set of important passages, $\mathcal{O}_{i}=\{o_{1},o_{2},\dots,o_{n}\}$. Each passage is important for generating a review for the paper it is associated with. This mapping is used as a training set for training the GNN. The node embeddings for sentence nodes are initialized with embeddings from a Sentence-Transformer (Reimers and Gurevych, 2019) model, namely all-mpnet-base-v2. The node embeddings for all other nodes are initialized with zero embeddings of dimension $768$, matching that of the Sentence-Transformer model.
Now, a Graph Attention Network (Veličković et al., 2018) is used to train the graph to predict the important passages. After passing the initial representation of the graph nodes through three layers of GATConv (PyTorch Geometric’s implementation of Graph Attention Networks) [^6] https://pytorch-geometric.readthedocs.io/en/latest/generated/torch_geometric.nn.conv.GATConv.html, we apply a non-linear classifier to the final passage node representations. The classification is defined as follows:
  1. (1)
    A label of 0 indicates that the passage is not important for review generation.
  2. (2)
    A label of 1 indicates that the passage is important for review generation.
The GNN is trained using various configurations based on the values of $k$ and $m$, as described in the DPR process in Section 4.1. These trained GNN models are denoted as GNN $(k,m)$.
We hypothesize that our choice of graph construction which models the semantic representation of nodes and flow of information through structure, allows our model to capture the intricate inter-dependencies of sentences and passages across different sections within an academic paper, and effectively make a choice about which passage(s) is/are important. The dependencies, according to our hypothesis, are learned from the sequential and hierarchical connections that we introduce in our graph construction method. The passing of information from one node to another is enhanced by the GATConv (Equations 2, 3, and 4) layers, which involves the concept of attention to selectively propagate messages in the Message Passing layer of a GNN.
(2) $\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$
(3) $e_{ij}=\text{LeakyReLU}\Bigl(\mathbf{a}^{T}\bigl[\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}\bigr]\Bigr)$
(4) $\alpha_{ij}=\frac{\exp\bigl(e_{ij}\bigr)}{\sum_{k\in\mathcal{N}(i)}\exp\bigl(e_{ik}\bigr)}$
Additional implementation details related to this section are available in Appendix B.
As illustrated in the example from our test set (Figure 3), the DPR-based approach fails to retrieve a key passage related to the main contributions of the paper, whereas our GNN successfully identifies it. Therefore, our multi-modal graph, which leverages both sequential and hierarchical node connections, more effectively identifies the salient passages required to enhance the review generation process. The GNN extracted passages, along with a structured prompt, serve as the input in the LLM fine-tuning stage, as discussed in Section 4.5.
4.4. Grounding Requirements and Inference Efficiency
For practical deployment, our GNN-based method offers a crucial advantage by operating autonomously at inference time. It identifies important passages directly from the paper’s graph structure, eliminating the need for grounding from human-generated reviews which act as queries in the DPR-based approaches, but may be unavailable in real-world scenarios.
Second, by selecting only the passages deemed most relevant, we significantly reduce the input length and associated computational costs for downstream LLM-based review generation. This not only leads to higher performance of LLMs, as outlined by Hsieh et al. (2024) but also leads to higher scores of the generated reviews compared to our baselines (see Section 5). The statistics related to the decrease in number of passages and token length post-extraction by the GNN model are reported in Table 2.
Table 2. Average token length (Avg. Token Len) and Average passage count (Avg. Passages) for a full paper, our best-configuration: GNN $(5,5)$, and our fastest-configuration: GNN $(5,3)$.
Input to LLM Avg Token Len Avg Passages
Full Paper 9445.92 47.07
GNN $(5,5)$ 4147.15 29.49
GNN $(5,3)$ 2899.59 21.00
4.5. Fine-Tuning Large Language Models for Review Generation
General-purpose LLMs are not trained specifically to generate peer reviews for academic research papers. Our motivation for fine-tuning is to ensure that the models learn to generate qualitatively sound, factual and structured peer reviews.
We select Mistral-Instruct-v0.2 (Mistral) [^7] https://arxiv.org/abs/2310.06825, inspired by our baseline SEA-EA, which also uses Mistral-Instruct-v0.2 as its underlying backbone model. We also use Llama 3.1 8B (Llama) [^8] https://arxiv.org/abs/2407.21783, as it is an open-weight state-of-the-art (SOTA) model with strong text generation capabilities for a variety of tasks. We employ LoRA (Low-Rank Adaptation) (Hu et al., 2022), for fine-tuning due to its efficiency in adapting large language models to particular tasks while maintaining computational feasibility. This allows us to leverage the pre-trained capabilities of Llama and Mistral while efficiently adapting them for review generation.
We use the trained GNN, as described in Section 4.3, to retrieve relevant passages for each paper that serve as the LLM’s input during fine-tuning. The human-written peer reviews for ICLR 2024 papers, which are consolidated with GPT-4, served as the ground truth for fine-tuning the LLM used to generate the final reviews. These reviews are released as a part of the dataset released by Yu et al. (2024). The dataset split is summarized in Table 3.
Table 3. Number of samples used for training and testing sourced from Yu et al. (2024)
Dataset Split Number of Samples
Training Set 1,765
Test Set 271
Both Mistral and Llama are fine-tuned on the passages retrieved, using various configurations of the trained GNN $(k,m)$, as described in Section 4.3. These fine-tuned models are denoted Mistral $(k,m)$ and Llama $(k,m)$, respectively. The prompt and hyperparameter settings used during fine-tuning are provided in Appendix C.
5. Results and Analysis
Table 4. Overall performance on the ICLR-24 reviews test dataset, comparing M-7B, SEA-E, SEA-EA as baselines from Yu et al. (2024) with our fine-tuned models. For both Llama-3.1-8B (Llama) and Mistral-Instruct-v0.2 (Mistral), the configuration with $(5,5)$ achieves the best performance, while $(5,3)$ has the shortest training time. Bold highlights the best scores, and underlined values indicate the second-best. Please note that the GPU configurations for finetuning Llama and Mistral were different. For details, refer to Appendix C.
Model ROUGE F1 ROUGE Recall BERTScore-F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
M-7B 22.55 8.89 11.28 13.48 5.29 16.73 83.79
SEA-E 34.29 11.71 14.98 23.28 7.90 10.13 84.04
SEA-EA 36.97 13.02 15.88 25.62 8.97 10.97 84.15
Mistral $(5,3)$ 53.44 15.47 22.39 53.18 15.42 22.31 86.44 4.34
Llama $(5,3)$ 54.24 15.91 22.59 53.75 15.77 22.41 86.59 3.60
Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
This section details our system’s performance analysis. Section 5.1 reports the evaluation of generated review quality using ROUGE-1, ROUGE-2, and ROUGE-L (Recall and F1) (Lin, 2004), as well as BERTScore-F1 (Zhang* et al., 2020), on the GPT-4-consolidated reviews test set for ICLR 2024 papers. ROUGE measures n-gram overlap, with recall reflecting content coverage and F1 balancing precision and recall, while BERTScore captures semantic similarity in embedding space. Section 5.2 further employs the Aspect Coverage (ACov) metric (Yuan et al., 2022), designed specifically for assessing peer-review quality.
Given the limitations of automated metrics, we also conduct human evaluations (Section 5.3) to assess review acceptability. To address the inherent scalability challenges of manual assessment, we employ an LLM-as-a-Judge framework (Section 5.4), which shows strong correlation with human judgments.
5.1. Performance Comparison with Baselines
We compare the performance of our models against baselines sourced from Yu et al. (2024). As reported in Table 4, the M-7B baseline performs direct inference with Mistral guided by an instruction prompt to generate peer reviews, SEA-E baseline leverages standardized fine-tuning data to produce more constructive reviews, while the SEA-EA baseline further introduces a mismatch score to evaluate the consistency between paper content and generated reviews. Our results suggest that we consistently outperform the baselines on all metrics, with significant improvements in the ROUGE metrics (Recall and F1). Our best-performing configuration, Llama $(5,5)$ (selected based on the highest BERTScore), achieves an improvement of 58.72% over the state-of-the-art (SOTA) baselines across all evaluation metrics. To ensure that these gains are not solely due to using a more capable LLM, we also evaluate our method using Mistral. The results demonstrate that even with Mistral, integrating our Multi Modal GNN-based passage selection method leads to substantial improvements, confirming that the performance boost is primarily due to our graph-based retrieval strategy rather than the choice of the LLM alone. The integration of a Graph Attention Network (GAT) enables the model to effectively capture complex dependencies between relevant passages and sentences within the paper, offering the essential context required to generate comprehensive reviews.
Our fastest, second-best, fine-tuned model, Llama $(5,3)$, performs competitively while being highly efficient. It reduces the training time by 32.07% with only a 0.71% drop in performance compared to our best model, consistently outperforming all baselines. This shows our GNN-based approach achieves near-optimal performance with significantly less computational overhead.
5.2. Aspect Coverage Analysis
As discussed by Yuan et al. (2022), each comment in a peer review typically addresses diverse aspects of an academic research paper, such as Summary (SUM), Motivation/Impact (MOT), Originality (ORI), Soundness/Correctness (SOU), Substance (SUB), Replicability (REP), Meaningful Comparison (CMP), and Clarity (CLA). These cannot be reliably assessed by conventional automatic metrics like ROUGE-F1 or BERTScore. To address this, Yuan et al. (2022) introduces an Aspect Tagger which can be used to compute Aspect Coverage (ACov) that measures the proportion of aspects covered by a peer-review. We evaluate the two best-performing models Llama $(5,5)$ and Llama $(5,3)$ using this tagger.
Table 5. Aspect Coverage (%) of model-generated reviews evaluated using the Aspect Tagger (Yuan et al., 2022). Summary (SUM), Clarity (CLA), Soundness/Correctness (SOU), Substance (SUB), Originality (ORI), Meaningful Comparison (CMP), Replicability (REP), and Motivation/Impact (MOT).
Model SUM CLA SOU SUB ORI CMP REP MOT
Llama (5,3) 100.0 99.6 98.9 95.2 92.6 79.0 72.0 66.8
Llama (5,5) 100.0 99.3 99.3 95.9 90.0 84.5 71.6 66.8
As shown in Table 5, both variants achieve strong coverage for Summary, Clarity, and Soundness, indicating the reliability of our model-generated reviews in capturing these core aspects. However, the aspect coverage for Motivation/Impact and Replicability is comparatively lower. Overall, the results suggest that our model captures most key aspects, indicating that the generated reviews are comprehensive and address the quality of multiple dimensions of a research paper.
5.3. Human Evaluation
Understanding the limitations of automatic evaluation metrics as highlighted by Hanna and Bojar (2021) and Schluter (2017), we conduct human evaluations to assess the acceptability of the generated reviews. We model our evaluation around two distinct aspects of review quality. The first is a Qualitative Assessment, where the generated review is evaluated in isolation across four key criteria: Reviewer Confidence, Review Thoroughness, Review Constructiveness, and Review Helpfulness. The second is a Factual Validity Assessment, which ensures the review’s specific claims are factually grounded in the source document. Since human evaluation is very costly, we conduct each of the two tasks on a subset of the test dataset, as described in their individual subsections. Additionally, we only considered Llama $(5,3)$ and Llama $(5,5)$, our best performing models for the evaluation process.
5.3.1. Qualitative Assessment
A total of 30 papers are chosen at random from our test set. Each of these papers are associated with 3 reviews, one Gold, one from Llama $(5,3)$, and one from Llama $(5,5)$, thus, resulting in a total set of 90 reviews. A total of 27 annotators are chosen, and each annotator is given 10 reviews to annotate. The evaluation criteria is designed to capture key aspects of review quality and include:
  1. (1)
    Reviewer Confidence (RCf): Reflects the reviews’ depth of technical understanding and engagement with the content.
  2. (2)
    Review Thoroughness (RT): Measures how well the review covers the methodology, results, and reproducibility.
  3. (3)
    Review Constructiveness (RCt): Assesses the clarity and usefulness of suggestions for improvement.
  4. (4)
    Review Helpfulness (RH): Indicates how well the review identifies important issues and explains them clearly to the authors.
Each annotator is tasked to assign a score from 1 to 5 for each of the above criteria. The evaluation guidelines are detailed in Appendix D. We end up with each review being annotated by 3 annotators. Score for a review is averaged across all its annotators, and the final score of a model, is the average score across all its reviews. The results of this experiment are summarized in Table 6.
Table 6. Results of our human qualitative assessment. This table shows average scores (on a scale of 1-5) comparing gold-standard reviews with those generated by our two best-performing Llama models across four key criteria: Reviewer Confidence (RCf), Review Thoroughness (RT), Review Constructiveness (RCt), and Review Helpfulness (RH).
Review RCf RT RCt RH
Gold 3.54 3.49 3.26 3.41
Llama (5,3) 3.37 3.30 3.27 3.34
Llama (5,5) 3.27 3.24 3.19 3.30
5.3.2. Factual Validity Assessment
A total of 10 papers are chosen at random from our test set. Each of these papers are associated with 3 reviews, one Gold, and two generated by our best performing models. Thus, the final evaluation set consists of 10 papers, each associated with 3 reviews, one Gold, one from Llama $(5,3)$, and one from Llama $(5,5)$. These 10 papers are circulated among 10 annotators. It is ensured that a single paper is annotated by at least 2 annotators. Annotators are instructed to evaluate each statement within a review based on its factual grounding in the full source paper. For each point, they assign a rating from a four-point scale: ‘Completely Agree’, ‘Mostly Agree’, ‘Mostly Disagree’, or ‘Completely Disagree’. The evaluation guidelines are detailed in Appendix D. The results of this experiment are summarized in Table 7 and Table 8.
Table 7. Detailed results of the Factual Validity Assessment. This table presents the percentage distribution of human agreement ratings for claims made in reviews from our Llama models and the gold standard. The abbreviations are: Completely Agree (CA), Mostly Agree (MA), Mostly Disagree (MD), and Completely Disagree (CD).
Review CA MA MD CD
Gold 39.12% 37.85% 16.40% 6.62%
Llama (5,3) 32.64% 31.94% 22.92% 12.50%
Llama (5,5) 32.27% 32.98% 23.05% 11.70%
Table 8. Detailed results of the Factual Validity Assessment. This table shows the overall Positive Rating Percentage, calculated by summing the “Completely Agree” and “Mostly Agree” categories from Table 7, which indicates the factual soundness of the reviews.
Review Positive Rating %
Gold 76.97%
Llama (5,3) 65.25%
Llama (5,5) 64.58%
5.4. LLM-as-a-Judge Evaluation
As mentioned by Zheng et al. (2023), the LLM-as-a-Judge method offers an explainable and affordable way to approximate human preferences, which are otherwise costly to gather. We employ a large language model, Gemini-2.0-Flash [^9] https://deepmind.google/technologies/gemini/ as an automated judge to evaluate the quality of reviews generated by our proposed model as compared to the gold-standard reviews. For the LLM-as-a-judge task, the evaluation criteria and prompting guidelines are adapted from those used in our human qualitative assessment, which are detailed in Section 5.3.1. Since LLM-as-a-Judge evaluation is scalable and much cheaper to undertake, we evaluate all our models against the gold review, over the entire test dataset. A key finding from our evaluation is the strong correlation between the LLM-as-a-Judge scores (Table 9) and the results from our human annotators (Table 6). This alignment suggests that the LLM-as-a-Judge is a reliable and valid proxy for assessing review quality in this domain.
Table 9. Qualitative evaluation of model-generated reviews using Gemini-2.0-Flash as the LLM judge. The table presents the average scores (out of 5) for all model configurations and the gold standard across four criteria: Reviewer Confidence (RCf), Review Thoroughness (RT), Review Constructiveness (RCt), and Review Helpfulness (RH). Bold values indicate the best score in each column (excluding Gold), and underlined values indicate the second-best.
Review RCf RT RCt RH
Gold 3.98 3.87 3.79 3.79
Mistral $(5,3)$ 3.50 3.35 3.44 3.32
Llama $(5,3)$ 3.61 3.45 3.47 3.39
Mistral $(5,5)$ 3.48 3.36 3.43 3.34
Llama $(5,5)$ 3.63 3.44 3.47 3.41
5.5. Ablation Studies
We perform ablation studies by systematically varying the values of $k\in\{1,3,5\}$ and $m\in\{3,5\}$ during LLM fine-tuning. The detailed results are reported in Appendix A. The results indicate that increasing $k$ and $m$ provides richer context to both Mistral and Llama to generate reviews, leading to consistent improvements across all evaluation metrics. Additionally, we observe that higher values of $m$ result in longer training times, as the GNN retrieves and passes a larger set of passages as input for finetuning the LLMs.
6. Limitations
While our framework provides a robust and high-performing solution for review generation, we acknowledge that no automatic system can fully replicate human judgment in complex tasks such as peer reviewing, as noted by Ye et al. (2024). Our goal is to assist authors in refining their papers and we aim for our system to support human reviewers. Currently, our system is primarily focused on ICLR 2024 papers due to the limited availability of public datasets containing consolidated reviews. However, we believe in the generalization capabilities of our model, given the proven effectiveness of graph-based approaches in various downstream NLP tasks, as discussed in Section 1.
Our reliance on human written reviews, consolidated using GPT-4 represents another limitation. While multi-review datasets more closely reflect the diversity of real peer review, they are often noisy and contradictory, which can hinder model learning. Consolidated reviews provide a clearer and more consistent supervision signal, but at the cost of reducing the richness of differing reviewer perspectives. Future work should explore hybrid approaches that retain both consistency and diversity.
7. Ethical Considerations
The deployment of AutoRev raises key ethical concerns related to bias, misuse, and academic integrity. Since the system is trained on existing peer reviews, it may inherit and amplify biases present in those texts, such as leniency toward well-known authors or under-representation of certain research areas, potentially skewing evaluations. Moreover, there is a risk that users may over-rely on AutoRev for making or substituting final review decisions, which could undermine the depth and critical thinking central to human peer review. Additionally, the realistic quality of AI-generated reviews opens the possibility of unethical use, such as submitting AI-authored critiques as original human work. To address these issues, we advocate for responsible usage of AutoRev as a supplementary tool, accompanied by transparency in AI involvement and safeguards to preserve academic integrity.
8. Conclusion
In this work, we introduce AutoRev, an automatic peer review generation system designed specifically for academic research papers. Our approach leverages a novel graph-based framework to represent the structure of a paper which enables the effective extraction of important passages using a Graph Neural Network. This strategy of selecting passages and reducing the input length to LLMs addresses the limitations of traditional methods, like finetuning using the entire paper as input (Yu et al., 2024), which leads to sub-optimal performance of LLMs due to large input token length (Hsieh et al., 2024).
Unlike DPR-based methods that require initial grounding on human written references to retrieve relevant passages, AutoRev employs a GNN that learns to identify important content directly from the paper’s graphical representation, without the need for grounding on existing reviews. This not only streamlines the process but also reduces computational overhead while enhancing review quality.
Our experiments on the ICLR 2024 dataset demonstrate that AutoRev consistently outperforms existing baselines, by 58.72% on popular evaluation metrics like ROUGE, and BERTScore. These findings are further strongly backed by human evaluations, where the generated reviews were rated nearly as constructive and acceptable as human-written ones, suggesting AutoRev could meaningfully reduce reviewer workload.
Beyond these performance gains, we emphasize AutoRev’s broader role in the web ecosystem. Academic research papers represent some of the most authoritative and trusted content on the web. By accelerating the peer-review workflow and enabling scalable, automated feedback, AutoRev contributes to the timely dissemination of credible scholarly knowledge online. In doing so, it supports not only authors and reviewers but also the wider web community that depends on accurate, trustworthy research for discovery, indexing, and downstream applications. Future work will explore extending this framework to a broader range of academic domains and investigating further improvements in graph construction and neural retrieval techniques to continue advancing the state-of-the-art in automated review generation.
Appendix A Ablation studies
The quantitative performance of different combinations of $k$ (number of review sentences concatenated; $k=1,3,5$) and $m$ (number of top DPR-retrieved passages used to train the GNN; $m=3,5$) for review generation is reported in Table 10 and Table 11, evaluated against the gold-standard reviews. While calculating the ROUGE metrics, we enable the use_stemmer option to ensure consistent matching of words with different morphological forms, to improve the reliability of the evaluation.
Table 10. Overall performance of Mistral across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
Model ROUGE-F1 ROUGE-Recall BERTScore F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
Mistral $(1,3)$ 53.54 15.81 22.68 53.26 15.74 22.59 86.51 7.60
Mistral $(1,5)$ 53.54 15.76 22.64 52.81 15.56 22.36 86.51 10.22
Mistral $(3,3)$ 53.86 15.80 22.62 53.21 15.63 22.36 86.54 6.33
Mistral $(3,5)$ 53.87 15.91 22.74 53.43 15.79 22.57 86.54 7.64
Mistral $(5,3)$ 53.44 15.47 22.39 53.18 15.42 22.31 86.44 4.34
Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
Table 11. Overall performance of Llama across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
Model ROUGE-F1 ROUGE-Recall BERTScore F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
Llama $(1,3)$ 54.42 16.19 22.76 53.71 15.98 22.48 86.61 6.65
Llama $(1,5)$ 54.40 16.09 22.82 53.58 15.86 22.50 86.61 8.23
Llama $(3,3)$ 54.52 16.03 22.67 53.79 15.83 22.39 86.61 5.11
Llama $(3,5)$ 54.61 16.30 22.88 53.78 16.06 22.56 86.66 6.28
Llama $(5,3)$ 54.24 15.91 22.59 53.75 15.77 22.41 86.59 3.60
Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
Appendix B GNN Implementation Details
We train and test our model on a single NVIDIA RTX6000 GPU with 48GB of VRAM.
B.1. Model Architecture
The model contains a Graph Neural Network (GNN) module and a passage classifier.
GNN Module
The GNN uses three graph attention (GATConv) layers. The first GATConv layer processes input features (of size 768 (embedding dimension of Sentence Encoder)) into 128-dimensional features with 4 attention heads, producing a concatenated output of $128\times 4$ dimensions. A residual connection is incorporated by applying a linear layer to project the input to $128\times 4$, which is added to the GATConv output. This is followed by an ELU activation and dropout ($p=0.2$).
The second GATConv layer maps the $128\times 4$ input to 256 dimensional features using 2 attention heads (output dimension $256\times 2$). A residual connection similarly projects the prior layer’s output to $256\times 2$, followed by ELU and dropout ($p=0.2$). The final GATConv layer upscales the $256\times 2$ features to the target output dimension (1024 in our case) without residual connections and dropout layers.
Passage Classifier
For passage nodes, we employ a feedforward classifier to map the output embeddings from the GNN to the final label space. The classifier is implemented as a multi-layer perceptron (MLP) comprising the following sequence of operations:
  • A linear transformation projecting the input to a 512 dimensional space.
  • A SiLU activation followed by dropout with a probability of $p=0.3$.
  • A second linear layer projecting to 256 dimensions.
  • Layer normalization applied to the 256-dimensional features.
  • Another SiLU activation.
  • A final linear layer projecting to the number of target classes (typically 2 for binary classification).
This architecture enables the model to learn nonlinear transformations of the graph-encoded passage representations while incorporating regularization and normalization for stable training.
B.2. Training
The model is trained for 30 epochs, with the best validation performance checkpoint retained for inference. The loss function used for training is the Cross Entropy Loss.
Appendix C Finetuning Implementation Details
We fine-tune Llama-3.1-8B using two NVIDIA L40S GPUs and Mistral using a single NVIDIA RTX 6000 Ada GPU. Both GPUs are equipped with 48GB of GDDR6 VRAM.
We apply Low-Rank Adaptation (LoRA) with a rank of 8 and a scaling factor ($\alpha=16$). LoRA is integrated into the following transformer layers: q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj. Models are fine-tuned for 3 epochs. The total training time for different configurations is reported in Table 10 and Table 11. The full set of hyperparameters used during training is listed in Table 12.
Table 12. Hyperparameters Used During Finetuning
Hyperparameter Value
Batch Size 1
Gradient Accumulation Steps 1
16-bit Floating Point Precision True
Optimizer adamw_torch
Learning Rate $1\times 10^{-4}$
Max Gradient Norm 0.3
Warmup Ratio 0.03
LoRA Rank 8
LoRA scaling factor 16
LoRA Dropout 0.05
Weight Decay 0.01
LR Scheduler Type cosine
C.1. Prompt for Finetuning
The prompt used for fine-tuning both Llama 3.1 8B and Mistral-Instruct-v0.2 is shown in Listing C.1.
[…]
The extracted relevant passages are inserted ahead of the ”Research Paper Passages” heading. Similarly, the corresponding content for each review section (Summary, Strengths, Weaknesses, and Questions) is added after their respective headings in the generated review.
Appendix D Human Evaluation
Due to space constraints, the detailed evaluation guidelines are made available in an anonymized GitHub repository: Anonymous Repository. For the camera-ready version, we will release these materials in a public GitHub repository to ensure full transparency and accessibility. We will now proceed to show the snapshots of our annotation UI, developed using Streamlit  [^10] https://streamlit.io/.
D.1. Qualitative Assessment
Examples of the user interface used for the Qualitative Assessment can be seen in Figure 4 and Figure 5.
Figure 4. This interface component corresponds to the Qualitative Assessment task described in Section 5.3.1. It marks the beginning of the review presentation. Due to space constraints, only the summary and a subset of the strengths from one review are shown here.:
Figure 5. This interface component corresponds to the task described in Section 5.3.1. In this part, annotators provide a final rating across the four categories and submit their annotation.:
D.2. Factual Validity Assessment
Examples of the user interface used for Factual Validity Assessment can be seen in Figure 6a and Figure 6b.
(a) The original paper is displayed in a PDF viewer to provide factual context.
(b) Each point of the review is rated for accuracy using a dropdown menu.
Figure 6. The user interface for the Factual Validity Assessment task described in Section 5.3.2. On the left (a), the full research paper is presented to the annotator. On the right (b), the annotator rates each point of the review based on its consistency with the source document.:
Appendix E Example of a Generated Review
We present a review generated by our system for the paper: On the Role of Edge Dependency in Graph Generative Models [^11] https://openreview.net/forum?id=LCQ7YTzgRQ
Summary
The paper introduces a novel framework for evaluating generative models of graphs, focusing on the concept of “overlap” to ensure both accuracy and edge diversity. It categorizes graph generative models into three levels of complexity: edge independent, node independent, and fully dependent models. The paper derives theoretical bounds on triangle counts and other short-length cycles based on overlap and provides new generative models for each level. However, the paper’s novelty and practical contributions are questioned due to the use of existing methods and the lack of significant empirical validation. The experimental section is seen as limited and does not convincingly demonstrate the superiority of the proposed methods over existing ones.
Strengths
  • The paper introduces a novel evaluation framework that considers the importance of model-generated graph overlap to ensure both accuracy and edge diversity, which is a significant contribution to the field of graph generative models.
  • The paper presents a hierarchy of graph generative models categorized into three levels of complexity, which encapsulates a wide range of prevalent methods and provides a structured framework for evaluating graph generative models.
  • The theoretical bounds on the number of triangles and other short-length cycles are well-presented and support the findings with instances demonstrating the asymptotic optimality of these bounds.
  • The authors introduce new generative models for each of the three hierarchical levels, leveraging dense subgraph discovery (Gionis & Tsourakakis, 2015), which is a practical application of the theoretical findings.
Weaknesses
  • The novelty of the paper is limited as it primarily relies on existing methods and theoretical frameworks, and the experimental results do not convincingly demonstrate the superiority of the proposed methods over existing ones.
  • The empirical evaluation is weak and does not provide sufficient evidence to support the claims made in the paper. The experimental results are not robust and the paper does not provide enough comparisons with other models to establish the effectiveness of the proposed methods.
  • The paper has several presentation issues, including unclear definitions and notations, which could hinder the understanding of the content for readers not familiar with graph generative models.
  • The practical applications and the significance of the proposed methods are not clearly demonstrated, and the paper lacks a detailed discussion on the limitations and potential future work.
Questions
  • Can the authors clarify the definitions and notations used in the paper, particularly the ones related to the overlap and triangle density, to ensure a clearer understanding of the content?
  • How do the authors justify the novelty of their work given the reliance on existing methods and theoretical frameworks?
  • What specific contributions does the paper make to the field of graph generative models, and can these be clearly articulated?
  • Why was the empirical evaluation conducted in a limited manner, and how can the paper be improved to provide more robust and convincing experimental results?
  • Could the authors discuss the practical applications of the proposed methods and the significance of their work in more detail?
  • Are there any plans to extend the theoretical bounds to higher-order cycles, and how might this affect the overall conclusions drawn in the paper?