### v3

# Graph-Guided Passage Retrieval for Author-Centric Structured Feedback

Abstract
Obtaining high-quality, pre-submission feedback is a critical bottleneck in the academic publication lifecycle for researchers. We introduce AutoRev, an automated author-centric feedback system that generates structured, actionable guidance prior to formal peer review. AutoRev employs a graph-based retrieval-augmented generation framework that models each paper as a hierarchical document graph, integrating textual and structural representations to retrieve salient content efficiently. By leveraging graph-based passage retrieval, AutoRev substantially reduces LLM input context length, leading to higher-quality feedback generation. Experimental results demonstrate that AutoRev significantly outperforms baselines across multiple automatic evaluation metrics, while achieving strong performance in human evaluations. Code will be released upon acceptance.
Graph-Guided Passage Retrieval for Author-Centric Structured Feedback
Maitreya Prafulla Chitale1thanks: Equal contribution.   Ketaki Mangesh Shetye [^1] footnotemark: 1 Harshit Gupta2Manav Chaudhary3Manish Shrivastava1Vasudeva Varma1 1IIIT Hyderabad  2Microsoft, India  3Sentisum {maitreya.chitale, ketaki.shetye}@research.iiit.ac.in guptaharshi@microsoft.com manav@sentisum.com {m.shrivastava, vv}@iiit.ac.in
1 Introduction
Recent advancements have focused on assisting reviewers and directly generating peer reviews Yu et al. (2024); Aitymbetov and Zorbas (2025); D’Arcy et al. (2024); Idahl and Ahmadi (2025). Concurrently, there has been growing interest in examining the strengths and limitations of such systems, particularly the use of large language models (LLMs) in this domain Ye et al. (2024). Relying on LLMs to generate or inform final review decisions remains risky due to issues of bias, overconfidence, and superficial feedback and these decision-oriented tools lack utility for authors in the pre-submission phase. Consequently, we advocate for a new paradigm towards author-centric support. We introduce AutoRev, an author-centric system that generates structured, actionable pre-submission feedback containing Summary, Strengths, Weaknesses, and Questions, without producing acceptance recommendations or replacing human peer review.
Figure 1: Example document graph with nodes for headings, sub-headings, passages, and sentences. This structure allows AutoRev to capture hierarchical and sequential dependencies across sections.
With the rise of Graph Neural Networks, the use of graphs as extractors or as guides for facilitating downstream tasks has recently gained traction in the NLP community Chitale et al. (2025); Gao et al. (2025); Agarwal et al. (2025); Zhu et al. (2025). Inspired by this, we leverage the graph’s capability as an extractor of passages to significantly reduce the input token length to the LLM. Reducing the token length of the input is shown to enhance LLM performance Hsieh et al. (2024). To summarize our main contributions:
  1. 1.
    We propose an author-centric system for generating concise, structured pre-submission feedback for academic papers.
  2. 2.
    We introduce a graph-based retrieval framework that learns to identify salient passages, reducing LLM input length while improving feedback quality.
  3. 3.
    We validate our approach using comprehensive human evaluations to assess feedback quality and factual grounding, in addition to standard automatic metrics.
2 Leveraging Peer Reviews as Proxy for Pre-Submission Feedback
In the absence of large-scale, publicly available datasets explicitly designed for pre-submission feedback, we use human-written peer reviews as a proxy, as they provide expert critiques of writing, methodology, and presentation aligned with the guidance authors seek prior to submission. In particular, the Weaknesses and Questions sections offer actionable critiques and clarifications. Since the dataset spans papers at varying stages of writing maturity, the associated reviews reflect feedback targeted at authors with differing levels of experience. We further support this choice through an impact-oriented user study (Section 6.3.4).
3 Related Works
Recent work has explored the use of NLP techniques to support scholarly writing and review workflows, including reviewer selection Aitymbetov and Zorbas (2025), meta-review generation Kumar et al. (2023), and automated review generation Yu et al. (2024). Many of these approaches rely on large language models, ranging from modular agent-based systems D’Arcy et al. (2024) to fine-tuned review generators Idahl and Ahmadi (2025). In parallel, a separate line of work focuses on author-assistive tools for specific writing sub-tasks, such as limitation suggestion Faizullah et al. (2024), citation recommendation and literature discovery Gu and Hahnloser (2023), and drafting individual sections like related work Hu and Wan (2014) or introductions Liebling et al. (2025). While these are effective within their scoped objectives, they are either reviewer-facing or narrowly targeted, and often operate on full-paper inputs at inference time. As a result, they raise concerns around hallucination, superficial feedback, computational cost, and overconfidence in model-generated judgments Ye et al. (2024), while failing to provide holistic, structured feedback that supports authors prior to submission.
Yu et al. (2024) introduce an automated paper reviewing framework (SEA) that standardizes reviews using GPT-4 OpenAI et al. (2024) consolidation and proposes a self-correction strategy during fine-tuning. We adopt SEA as our baseline as it is trained on the publicly available ICLR 2024 dataset and generates the same structured fields as AutoRev, among others.
Studies have explored graph-based extensions to Retrieval-Augmented Generation (RAG) Lewis et al. (2020) to improve grounding and reasoning in LLMs. Zhu et al. (2025) provide a comprehensive taxonomy of Graph-RAG systems based on database construction, retrieval mechanisms, and processing pipelines. Within this framework, AutoRev represents a domain-specific Graph-RAG approach that constructs a hierarchical graph directly from a scientific paper and employs a learning-based Graph Attention Network veličković2018graph for passage retrieval.
4 Dataset
The dataset used in our study is borrowed from Yu et al. (2024), comprising papers submitted to various conferences along with multiple peer reviews per paper. Among these, the ICLR 2024 and NeurIPS 2023 datasets are publicly released, though the GPT-4-consolidated reviews for NeurIPS 2023 are not available. Hence, our analysis focuses only on the ICLR 2024 papers. Table 1 summarizes statistics for the ICLR 2024 papers Yu et al. (2024).
Attribute Value
Total Papers 5,653
Average Tokens per Paper 9,815
Total Reviews 21,839
Acceptance Rate 37%
Domain Machine Learning
Table 1: Dataset statistics for the ICLR 2024 dataset, as reported in Yu et al. (2024).
Each paper is associated with multiple reviews containing sections such as Summary, Strengths, Weaknesses, Questions, Soundness, Presentation, Contribution, Confidence, and Rating. We retain only the Summary, Strengths, Weaknesses, and Questions sections, as these encode constructive feedback for authors-centric pre-submission support (see Section 2), and exclude rating-based fields that correspond to final decision-making. For training, we use the GPT-4 consolidated reviews released by Yu et al. (2024), which merge multiple human-written reviews into a single structured review per paper.
To assess cross-venue performance, we curate a supplementary dataset of 10 papers from COLM and NeurIPS 2025 (CNT_10). This set complements the ICLR 2024 test data (ICT) for the human evaluation detailed in Section 6.3.
Figure 2: The system begins with a parsed document from which DPR extracts passages for GNN training. The GNN learns to identify key passages relevant to structured feedback generation. The trained GNN selects important passages without requiring human-generated feedback. These passages, combined with a structured prompt, are input to fine-tune an LLM. The fine-tuned LLM generates the final structured feedback, without external grounding.
5 Methodology
In this section, we describe the implementation of our proposed framework. Section 5.1 describes the initial passage extraction using Dense Passage Retrieval (DPR) (Karpukhin et al., 2020). The extracted passages are used as training data in Section 5.3, to train a Graph Neural Network (GNN) that learns to extract relevant passages from a research paper. The trained GNN enables our model to perform competitively on the test set without the need for grounding on human-generated feedback, as discussed in Section 5.4. Post that, we describe our process of fine-tuning LLMs to generate the final structured feedback, in Section 5.5. An overview of our methodology can be found in Figure 2.
5.1 DPR-based Passage Extraction
Dense Passage Retrieval (DPR) is a neural retrieval model trained primarily on question-answering datasets that encodes queries and passages into dense vector representations. Given an input query, DPR retrieves relevant passages based on semantic similarity in the embedding space. We use a pre-trained DPR model to identify passages corresponding to the Summary, Strengths, Weaknesses, and Questions sections of reviews for each paper in the training set. Specifically, we use the context encoder [^1] https://huggingface.co/facebook/dpr-ctx_encoder-multiset-base and query encoder [^2] https://huggingface.co/facebook/dpr-question_encoder-multiset-base from the multiset-base DPR variant.
[…]
This final set of passages represents the most relevant excerpts from the full paper that are semantically linked to the feedback content. These extracted passages serve as the ground truth labels for training the GNN, as described in Section 5.3
5.2 Graph Construction for Representing a Paper
The idea of graph construction employed in our work is inspired by Zhao et al. (2024). We borrow their idea of representing a document as a hierarchical graph but extend it beyond sentence level. The graph is constructed directly from the parsed paper and is designed to generalize across diverse academic writing styles and section hierarchies.
Formally, each paper is represented as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where nodes correspond to textual units at different granularities, including the paper root, section headings, sub-headings, passages, and sentences. Edges encode two complementary relationships: (i) hierarchical links ($\mathcal{E}_{hier}$) that reflect the document’s structural nesting and (ii) sequential links ($\mathcal{E}_{seq}$) that preserve the linear ordering of passages and sentences. Together, these connections allow information to propagate both vertically across document structure and laterally across adjacent text. A schematic illustration of the document graph is shown in Figure 1. Full details of node types, edge definitions, and construction procedures are provided in Appendix A.
This design explicitly mitigates over-squashing in tree-like graphs, where exponential neighborhood growth leads to information bottlenecks at the root (Alon and Yahav, 2021; Topping et al., 2022). Following the rewiring strategies of Topping et al. (2022), we augment the hierarchical backbone with sequential edges to facilitate lateral information flow and reduce reliance on hierarchical connections. Additionally, the inherent structure of scientific papers results in graphs of shallow depth (typically $\leq 5$ hops), limiting long-range dependency compression and preserving fine-grained information from leaf nodes.
5.3 GNN Training for Passage Retrieval
To train a GNN on the graph constructed in Section 5.2, we take the data generated in Section 5.1, which maps each paper $p_{i}$ to a set of important passages, $\mathcal{O}_{i}=\{o_{1},o_{2},\dots,o_{n}\}$. Each passage is important for generating a feedback for the paper it is associated with. This mapping is used as a training set for training the GNN. The node embeddings for sentence nodes are initialized with embeddings from a Sentence-Transformer Reimers and Gurevych (2019) model, namely all-mpnet-base-v2. The node embeddings for all other nodes are initialized with zero embeddings of dimension $768$, matching that of the Sentence-Transformer model.
We train a Graph Attention Network (GAT) veličković2018graph over the document graph to classify passage nodes as feedback-relevant or not. The GNN is trained using various configurations based on the values of $k$ and $m$, as described in the DPR process in Section 5.1. These trained GNN models are denoted as GNN $(k,m)$. Additional implementation details related to this section are available in Appendix C.
As illustrated in the example from our test set (Figure 3), the DPR-based approach fails to retrieve a key passage related to the main contributions of the paper, whereas our GNN successfully identifies it. Therefore, graph-based representation, which leverages both sequential and hierarchical node connections, effectively identifies the salient passages required to enhance the feedback generation process. The GNN extracted passages, along with a structured prompt, serve as the input in the LLM fine-tuning stage, as discussed in Section 5.5.
Figure 3: Excerpt of a test sample where passages highlighted in cyan colour refer to relevant passages identified by GNN only and yellow colour refer to relevant passages identified by both GNN and DPR
5.4 Grounding and Inference Efficiency
For practical deployment, our GNN-based method offers a crucial advantage by operating autonomously at inference time. Unlike DPR-based approaches, which rely on human-generated feedback to act as queries, that may be unavailable in real-world scenarios, our system identifies salient passages directly from the document graph, thereby eliminating the dependency on external grounding.
Unlike SEA baselines that process the full text, AutoRev generates feedback from a reduced set of GNN-selected passages. This approach significantly lowers computational cost and input length. As noted by Hsieh et al. (2024), shorter contexts enhance LLM performance, which is also reflected in our higher scores compared to baselines (Section 6). Table 2 details the specific reductions in tokens and passages.
Method Avg Token Len Avg Passages
Full Paper 9445.92 47.07
GNN $(5,5)$ 4147.15 29.49
GNN $(5,3)$ 2899.59 21.00
Table 2: Average token length (Avg Token Len) and Average passage count (Avg. Passages) for a full paper, our best-configuration: GNN $(5,5)$, and our fastest-configuration: GNN $(5,3)$.
5.5 Fine-Tuning Large Language Models for Structured Feedback Generation
General-purpose LLMs are not trained specifically to generate structured feedback for academic research papers. Our motivation for fine-tuning is to ensure that the models learn to generate qualitatively sound, factual and structured feedback.
We select Mistral-Instruct-v0.2 (Mistral) Jiang et al. (2023) to align with our SEA-EA baseline, and state-of-the-art (SOTA), open-weight model, Llama 3.1 8B (Llama) Grattafiori et al. (2024) for its strong text generation capabilities. To efficiently adapt these models for structured feedback generation while ensuring computational feasibility, we employ Low-Rank Adaptation (LoRA) Hu et al. (2022) for finetuning.
We use the trained GNN, as described in Section 5.3, to retrieve relevant passages from each paper, which are provided as input to the LLM during fine-tuning. For supervision, we use the Summary, Strengths, Weaknesses, and Questions sections of human-written peer reviews for ICLR 2024 papers, consolidated using GPT-4 and released as part of the dataset introduced by Yu et al. (2024). These consolidated reviews serve as structured feedback targets for training the LLM. The dataset split is summarized in Table 3.
Dataset Split Number of Samples
Training Set 1,765
Test Set 271
Table 3: Number of samples used for training and testing sourced from Yu et al. (2024)
Both models are fine-tuned on passages retrieved via varying GNN $(k,m)$ configurations (Section 5.3), denoted as Mistral $(k,m)$ and Llama $(k,m)$. Prompts and hyperparameters are detailed in Appendix D.
6 Results and Analysis
Model ROUGE F1 ROUGE Recall BERTScore-F1 Training Hours
R-1 R-2 R-L R-1 R-2 R-L
M-7B 22.55 8.89 11.28 13.48 5.29 16.73 83.79
SEA-E 34.29 11.71 14.98 23.28 7.90 10.13 84.04
SEA-EA 36.97 13.02 15.88 25.62 8.97 10.97 84.15
Mistral $(5,3)$ 53.44 15.47 22.39 53.18 15.42 22.31 86.44 4.34
Llama $(5,3)$ 54.24 15.91 22.59 53.75 15.77 22.41 86.59 3.60
Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
Table 4: Performance on the ICLR-24 test set comparing our models with baselines (M-7B, SEA-E, SEA-EA) from Yu et al. (2024). Bold and underlined values indicate best and second-best scores, respectively. Please note that the GPU configurations differ between Llama and Mistral fine-tuning (details in Appendix D).
This section presents a comprehensive evaluation of AutoRev across automatic, human, and robustness-focused analyses. We first benchmark feedback quality on the ICLR 2024 test set using ROUGE-1/2/L (Recall and F1) Lin (2004) and BERTScore-F1 Zhang* et al. (2020) against the sections in the reference feedback, assessing both lexical overlap and semantic similarity. To evaluate feedback quality beyond surface similarity, we additionally report Aspect Coverage (ACov) Yuan et al. (2022), which measures the breadth of scholarly aspects addressed in the generated feedback. Given the limitations of automated metrics, we conduct human evaluations assessing qualitative dimensions such as feedback depth, constructiveness, thoroughness, and helpfulness, alongside a distinct assessment of factual validity of the generated feedback. We further examine cross-venue generalization on NeurIPS and COLM 2025 papers and assess practical utility via an impact-oriented user survey. Finally, we employ a scalable LLM-as-a-Judge framework and perform ablation studies to analyze the effects of passage-selection parameters.
6.1 Performance Comparison with Baselines
We compare AutoRev against SEA-based baselines from Yu et al. (2024), including M-7B, SEA-E, and SEA-EA. All baselines operate on the full paper text during both training and inference, without any learned passage selection. M-7B performs direct inference using an instruction-tuned Mistral model, SEA-E leverages standardized fine-tuning to improve review quality, and SEA-EA further incorporates a mismatch score to assess consistency between paper content and generated review.
In contrast, AutoRev employs learned passage selection using a Graph Neural Network over a document graph, providing only salient passages to the LLM. As shown in Table 4 , AutoRev consistently outperforms all baselines across ROUGE (Recall and F1) and BERTScore metrics. Our best-performing configuration, Llama $(5,5)$ , achieves an average improvement of 58.72% over the state-of-the-art baselines.
To isolate the effects of model capacity from retrieval quality, we also evaluate AutoRev using Mistral. The substantial gains observed with Mistral confirm that the performance improvements stem primarily from GNN-based passage selection rather than the choice of LLM. The graph-based retrieval enables the model to capture both hierarchical and sequential dependencies across the document, providing the focused context necessary for generating comprehensive feedback.
Finally, Llama $(5,3)$ offers a strong efficiency-performance trade-off, reducing training time by 32.07% with only a 0.71% decrease in performance relative to Llama $(5,5)$ , while still outperforming all baselines.
6.2 Aspect Coverage Analysis
We evaluate the breadth of scholarly criteria addressed in the generated structured feedback using Aspect Coverage (ACov), following the aspect tagging framework of Yuan et al. (2022). Unlike surface-level similarity metrics, ACov measures whether feedback engages with the diverse dimensions typically considered in peer review.
Table 5 reports aspect coverage for our two best-performing configurations, Llama $(5,3)$ and Llama $(5,5)$. Both variants demonstrate consistently strong coverage across core feedback dimensions, except for Motivation/Impact and Replicability, indicating that AutoRev generates feedback that goes beyond summarization and engages with substantive and technical aspects of the paper.
Overall, these results suggest that AutoRev provides broad and balanced coverage of evaluation criteria, reinforcing its suitability as a comprehensive pre-submission feedback system.
Model SUM CLA SOU SUB ORI CMP REP MOT
Llama $(5,3)$ 100.0 99.6 98.9 95.2 92.6 79.0 72.0 66.8
Llama $(5,5)$ 100.0 99.3 99.3 95.9 90.0 84.5 71.6 66.8
Table 5: Aspect Coverage (%) of model-generated structured feedback evaluated using the Aspect Tagger (Yuan et al., 2022) across aspects: Summary (SUM), Clarity (CLA), Soundness/Correctness (SOU), Substance (SUB), Originality (ORI), Meaningful Comparison (CMP), Replicability (REP), and Motivation/Impact (MOT).
6.3 Human Evaluation
Understanding the limitations of automatic evaluation metrics, as highlighted by Hanna and Bojar (2021) and Schluter (2017), we conduct human evaluations to assess the acceptability, factual grounding, and practical utility of the generated feedback. Our evaluation is structured around two complementary dimensions: (i) qualitative assessment of feedback quality and (ii) factual validity of individual claims. We perform these evaluations on both the ICLR 2024 test set and a small out-of-domain dataset drawn from NeurIPS and COLM 2025. Due to the cost of expert annotation, human evaluations are conducted on representative subsets of the data. We report results for Llama $(5,3)$ and Llama $(5,5)$, our best-performing configurations.
6.3.1 Qualitative Assessment
A total of 30 papers are randomly selected from the ICLR 2024 test set (ICT_30). Each paper is associated with three feedback instances: one reference feedback corresponding to the ICT_30 paper and one generated by each of Llama $(5,3)$ and Llama $(5,5)$, yielding 90 feedback samples. A total of 27 annotators participate in the study, with each feedback annotated by three independent annotators. Feedback quality is evaluated across four criteria:
  1. 1.
    Feedback Depth (FD): Evaluates the technical understanding and meaningful engagement with the paper’s content.
  2. 2.
    Feedback Thoroughness (FT): Measures the coverage of key components like methodology, results, and reproducibility.
  3. 3.
    Feedback Constructiveness (FC): Assesses the utility of suggestions for improvement.
  4. 4.
    Feedback Helpfulness (FH): Evaluates how effectively important issues are identified and explained to the authors.
Annotators assign scores on a 1-5 scale following the guidelines detailed in Appendix E and snapshots of the annotation interface are shown in Appendix F. Scores are averaged across annotators for each feedback and then aggregated across all samples for each model.
As shown in Table 6, AutoRev-generated feedback closely approaches ICT_30 feedback in constructiveness and helpfulness, which are particularly critical for pre-submission support. While ICT_30 feedback remains superior in depth and thoroughness, the results show that AutoRev does not merely summarize paper content, but meaningfully engages with methodological and clarity issues, reinforcing its suitability as a structured feedback tool for authors.
Feedback FD FT FC FH
ICT_30 3.54 3.49 3.26 3.41
Llama $(5,3)$ 3.37 3.30 3.27 3.34
Llama $(5,5)$ 3.27 3.24 3.19 3.30
Table 6: Human Qualitative Assessment scores (scale of 1–5) of generated feedback using Llama models against reference feedback corresponding to the ICT_30 paper across four criteria: Feedback Depth (FD), Feedback Thoroughness (FT), Feedback Constructiveness (FC), and Feedback Helpfulness (FH).
6.3.2 Factual Validity Assessment
To evaluate factual grounding, we randomly select 10 papers from the ICLR 2024 test set (ICT_10). Each paper is associated with three feedback instances: one reference feedback corresponding to the ICT_10 paper and outputs from Llama $(5,3)$ and Llama $(5,5)$. These feedbacks are annotated by 10 annotators, with each paper evaluated by two annotators.
Annotators assess each statement within a feedback based on its factual consistency with the source paper, using a four-point scale: Completely Agree, Mostly Agree, Mostly Disagree, and Completely Disagree. The evaluation guidelines are provided in Appendix G, and snapshots of the annotation interface are shown in Appendix F.
The percentage agreement scores summarized in Tables 7 and 12 indicate that AutoRev maintains strong factual grounding in the source document. While ICT_10 feedback achieves the highest agreement, AutoRev-generated feedback receives a clear majority of positive ratings, with complete disagreements (potential hallucinations) remaining below 15%. These findings suggest that graph-based passage selection effectively anchors feedback generation in relevant evidence, mitigating hallucinated content.
6.3.3 Cross-Venue Generalization Assessment
To further examine robustness beyond the training distribution, we conduct an out-of-domain human evaluation on a curated set of 10 papers, comprising five papers each from COLM 2025 and NeurIPS 2025 (CNT_10). Each paper is evaluated using feedback generated by Llama $(5,3)$ and Llama $(5,5)$. Due to the absence of publicly available consolidated reference feedback for these venues and the high cost of expert annotation, reference feedback is not included.
A total of 10 annotators participate in the study, with each paper evaluated by two annotators following the same guidelines described in Section 6.3.2. The results, reported in Table 7 and 12, show comparable factual agreement across venues. This suggests that AutoRev generalizes beyond the ICLR 2024 training distribution, capturing structural and semantic regularities common to scientific writing rather than overfitting to venue-specific conventions.
Dataset Feedback CA MA MD CD
ICT_10 Reference 39.12 37.85 16.40 6.62
Llama $(5,3)$ 32.64 31.94 22.92 12.50
Llama $(5,5)$ 32.27 32.98 23.05 11.70
CNT_10 Llama $(5,3)$ 22.00 39.86 24.57 13.57
Llama $(5,5)$ 24.06 40.25 22.19 13.50
Table 7: Factual validity assessment (% agreement) on ICLR 2024 (ICT_10) and cross-venue CNT_10 datasets. CA/MA indicate grounded statements, while CD reflects clear factual errors (hallucinations).
6.3.4 Impact Assessment
To evaluate the impact of our system, we conduct a survey among the annotators involved in the tasks discussed in Sections 6.3.2 and 6.3.3. Participants are asked: “As an author, how likely are you to use AutoRev to get feedback for your own manuscript before submitting it to a conference?” On a 5-point scale, the system received an average rating of 4.2. This indicates that authors are strongly inclined to use our model for pre-submission critique, demonstrating the system’s practical utility.
6.4 LLM-as-a-Judge Evaluation
As mentioned by Zheng et al. (2023), the LLM-as-a-Judge method offers an explainable and affordable way to approximate human preferences, which are otherwise costly to gather. We employ a large language model, Gemini-2.0-Flash [^3] https://deepmind.google/technologies/gemini/ as an automated judge to evaluate generated feedback. Since LLM-as-a-Judge evaluation is scalable and much cheaper to undertake, we evaluate all our models against the reference feedback, over the entire ICT dataset. The evaluation criteria and prompting guidelines are adapted from those used in our human qualitative assessment (Section 6.3.1; Appendix E). The results of this experiment are summarized in Table 8
A key finding is the strong correlation between LLM-as-a-Judge scores and human annotations (Table 6). This alignment validates the automated judge as a reliable proxy for assessing feedback quality in this domain.
Feedback FD FT FC FH
ICT 3.98 3.87 3.79 3.79
Mistral $(5,3)$ 3.50 3.35 3.44 3.32
Llama $(5,3)$ 3.61 3.45 3.47 3.39
Mistral $(5,5)$ 3.48 3.36 3.43 3.34
Llama $(5,5)$ 3.63 3.44 3.47 3.41
Table 8: LLM-as-a-Judge (Gemini-2.0-Flash) scores (scale of 1–5) comparing model-generated feedback against ICT across Feedback Depth (FD), Feedback Thoroughness (FT), Feedback Constructiveness (FC), and Feedback Helpfulness (FH). Bold and underlined values indicate the best and second-best model scores (excluding reference feedback), respectively.
6.5 Ablation Studies
We perform ablation studies by systematically varying the values of $k\in\{1,3,5\}$ and $m\in\{3,5\}$ during LLM fine-tuning. The detailed results are reported in Appendix B. The results indicate that increasing $k$ and $m$ provides richer context to both Mistral and Llama to generate feedback, leading to consistent improvements across all evaluation metrics. Though, this comes at the downside of higher training times as $m$ increases for a fixed $k$.
7 Conclusion
In this work, we present AutoRev, an author-centric system for generating structured pre-submission feedback using graph-based passage retrieval. By modeling papers as hierarchical document graphs and leveraging a GNN to identify salient content, AutoRev substantially reduces LLM input length, overcoming the limitations of full-paper fine-tuning, stemming from long-context degradation Hsieh et al. (2024), while operating without human-written grounding. Experiments on ICLR 2024 show that AutoRev demonstrates strong empirical performance through automatic metrics, with human evaluations confirming constructive, well-grounded feedback and cross-venue generalization. Notably, Llama $(5,5)$ excels in generating comprehensive feedback for rigorous refinement of manuscripts, whereas Llama $(5,3)$ offers a favorable efficiency-performance trade-off for large-scale deployment.
Limitations
While our framework provides a robust and high-performing solution for feedback generation, we acknowledge that no automatic system can fully replicate human judgment in complex tasks such as peer reviewing, as noted by Ye et al. (2024). Our goal is to assist authors in refining their papers and we aim for our system to support authors only in the pre-submission feedback phase. Currently, our system is primarily focused on ICLR 2024 papers due to the limited availability of public datasets containing consolidated reviews. However, the consistent positive performance observed in out-of-domain human evaluations (Section 6.3.3) suggests that our model generalizes well beyond the ICLR 2024 setting.
Another limitation is our reliance on human-written reviews consolidated using GPT-4. While consolidated reviews provide a clearer and more consistent supervision signal than multi-review datasets, which are often noisy and contradictory and can hinder model learning, this approach introduces its own challenges. In particular, relying on an LLM for consolidation may induce biases. Future work can explore hybrid strategies that reduce or eliminate reliance on LLMs for consolidating multiple reviews into a single supervision signal.
Ethical Considerations
The deployment of AutoRev raises ethical concerns regarding bias, misuse, and academic integrity. Critically, while the system is designed strictly as an author-centric aid, there is a risk of over-reliance or misuse in decision-making, which could undermine the nuances of human reviewing. Furthermore, by training on historical peer reviews and GPT-4 consolidated data, the model may inherit and amplify existing biases present in those texts, such as leniency toward well-known authors or under-representation of certain research areas, potentially skewing evaluations. Additionally, the high fidelity of the generated text raises concerns about academic misconduct, where AI critiques could be misrepresented as original human work. To address these issues, we advocate for transparency and emphasize that AutoRev is intended solely as a supplementary tool to support, not replace, the scientific review process.
Appendix A Graph Construction
Formally, each paper is represented as an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where nodes $\mathcal{V}$ represent text segments with semantically encoded features, and $\mathcal{E}$ denote the hierarchical and sequential relationships that form the document’s structure. Our graph construction is designed to be generalizable across diverse academic paper structures, accommodating arbitrary levels of nesting and heterogeneous section organization.
A.1 Node Definitions
  • Paper Node ($v_{\text{paper}}$): A single root node representing the entire document.
  • Heading Nodes ($\mathcal{H}=\{h_{1},h_{2},\dots,h_{m}\}$): Nodes corresponding to top-level section headings.
  • Sub-heading Nodes ($\mathcal{S}_{i}=\{s_{i1},s_{i2},\dots,s_{ik}\}$): Nodes representing sub-headings under each heading $h_{i}$. This structure naturally extends to accommodate multiple levels of nesting, making it suitable for documents with deeper hierarchical organization.
[…]
  • Sentence Nodes ($\mathcal{T}_{jk}=\{t_{jk1},t_{jk2},\dots,t_{jkp}\}$): Nodes for individual sentences within each passage $o_{jk}$, tokenized using NLTK Bird and Loper (2004).
A.2 Edge Definitions
Edges capture both hierarchical and sequential relationships within the document.
  • Root Connections: $(v_{\text{paper}},h_{i})\in\mathcal{E}_{\text{hier}}$ for all $h_{i}\in\mathcal{H}$.
  • Heading-SubHeading Connections:
    $(h_{i},s_{ij})\in\mathcal{E}_{\text{hier}}$ for all $s_{ij}\in\mathcal{S}_{i}$.
  • Content Links:
    $(s_{ij},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ if }\mathcal{S}_{i}\neq\emptyset;(h_{i},o_{jk})\in\mathcal{E}_{\text{hier}}\text{ otherwise.}$
  • Passage-Sentence Connections:
    $(o_{jk},t_{jkq})\in\mathcal{E}_{\text{hier}}$ for all $t_{jkq}\in\mathcal{T}_{jk}$.
  • Sequential Edges:
    For passages: $(o_{jk},o_{j(k+1)})\in\mathcal{E}_{\text{seq}}$ for $1\leq k<n$.
    For sentences: $(t_{jkq},t_{jk(q+1)})\in\mathcal{E}_{\text{seq}}$ for $1\leq q<p$.
The final edge set is the union of hierarchical and sequential edges:
$\mathcal{E}=\mathcal{E}_{\text{hier}}\cup\mathcal{E}_{\text{seq}},$ (1)
where $\mathcal{E}_{\text{hier}}$ captures the vertical (structural) relationships and $\mathcal{E}_{\text{seq}}$ captures the horizontal (textual progression) flow. This graph construction is agnostic to section depth and flexibly adapts to any nested structure within academic documents. A toy example of such a graph, built from a paper with 2 Headings, 2 Sub-Headings, 4 Passages, and 6 Sentences, is shown in Figure 1.
Appendix B Ablation studies
The quantitative performance of different combinations of $k$ (number of review sentences concatenated; $k=1,3,5$) and $m$ (number of top DPR-retrieved passages used to train the GNN; $m=3,5$) for review generation is reported in Table 9 and Table 10, evaluated against the reference feedback. While calculating the ROUGE metrics, we enable the use_stemmer option to ensure consistent matching of words with different morphological forms, to improve the reliability of the evaluation.
Model ROUGE-F1 ROUGE-Recall BERTScore F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
Mistral $(1,3)$ 53.54 15.81 22.68 53.26 15.74 22.59 86.51 7.60
Mistral $(1,5)$ 53.54 15.76 22.64 52.81 15.56 22.36 86.51 10.22
Mistral $(3,3)$ 53.86 15.80 22.62 53.21 15.63 22.36 86.54 6.33
Mistral $(3,5)$ 53.87 15.91 22.74 53.43 15.79 22.57 86.54 7.64
Mistral $(5,3)$ 53.44 15.47 22.39 53.18 15.42 22.31 86.44 4.34
Mistral $(5,5)$ 53.65 15.80 22.74 53.60 15.79 22.74 86.55 6.44
Table 9: Overall performance of Mistral across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
Model ROUGE-F1 ROUGE-Recall BERTScore F1 Training Time (hrs)
R-1 R-2 R-L R-1 R-2 R-L
Llama $(1,3)$ 54.42 16.19 22.76 53.71 15.98 22.48 86.61 6.65
Llama $(1,5)$ 54.40 16.09 22.82 53.58 15.86 22.50 86.61 8.23
Llama $(3,3)$ 54.52 16.03 22.67 53.79 15.83 22.39 86.61 5.11
Llama $(3,5)$ 54.61 16.30 22.88 53.78 16.06 22.56 86.66 6.28
Llama $(5,3)$ 54.24 15.91 22.59 53.75 15.77 22.41 86.59 3.60
Llama $(5,5)$ 54.47 16.26 22.87 53.49 15.98 22.48 86.68 5.30
Table 10: Overall performance of Llama across different $(k,m)$ settings. Metrics include ROUGE-F1 (R-1, R-2, R-L), ROUGE-Recall (R-1, R-2, R-L), BERTScore F1 and training time (in hours). Bold highlights the best scores, and underlined values indicate the second-best.
Appendix C GNN Implementation Details
The passing of information from one node to another is enhanced by the GATConv (Equations 2, 3, and 4) layers, which involves the concept of attention to selectively propagate messages in the Message Passing layer of a GNN.
$\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}(i)}\alpha_{ij}\mathbf{W}\mathbf{h}_{j}\right)$ (2)
$e_{ij}=\text{LeakyReLU}\Bigl(\mathbf{a}^{T}\bigl[\mathbf{W}\mathbf{h}_{i}\parallel\mathbf{W}\mathbf{h}_{j}\bigr]\Bigr)$ (3)
$\alpha_{ij}=\frac{\exp\bigl(e_{ij}\bigr)}{\sum_{k\in\mathcal{N}(i)}\exp\bigl(e_{ik}\bigr)}$ (4)
We train and test our model on a single NVIDIA RTX6000 GPU with 48GB of VRAM.
C.1 Model Architecture
The model contains a Graph Neural Network (GNN) module and a passage classifier.
GNN Module
The GNN uses three graph attention (GATConv) layers. The first GATConv layer processes input features (of size 768 (embedding dimension of Sentence Encoder)) into 128-dimensional features with 4 attention heads, producing a concatenated output of $128\times 4$ dimensions. A residual connection is incorporated by applying a linear layer to project the input to $128\times 4$, which is added to the GATConv output. This is followed by an ELU activation and dropout ($p=0.2$).
The second GATConv layer maps the $128\times 4$ input to 256 dimensional features using 2 attention heads (output dimension $256\times 2$). A residual connection similarly projects the prior layer’s output to $256\times 2$, followed by ELU and dropout ($p=0.2$). The final GATConv layer upscales the $256\times 2$ features to the target output dimension (1024 in our case) without residual connections and dropout layers.
Passage Classifier
For passage nodes, we employ a feedforward classifier to map the output embeddings from the GNN to the final label space. The classifier is implemented as a multi-layer perceptron (MLP) comprising the following sequence of operations:
  • A linear transformation projecting the input to a 512 dimensional space.
  • A SiLU activation followed by dropout with a probability of $p=0.3$.
  • A second linear layer projecting to 256 dimensions.
  • Layer normalization applied to the 256-dimensional features.
  • Another SiLU activation.
  • A final linear layer projecting to the number of target classes (typically 2 for binary classification).
This architecture enables the model to learn nonlinear transformations of the graph-encoded passage representations while incorporating regularization and normalization for stable training.
C.2 Training
The model is trained for 30 epochs, with the best validation performance checkpoint retained for inference. The loss function used for training is the Cross Entropy Loss.
Appendix D Finetuning Implementation Details
We fine-tune Llama-3.1-8B using two NVIDIA L40S GPUs and Mistral using a single NVIDIA RTX 6000 Ada GPU. Both GPUs are equipped with 48GB of GDDR6 VRAM.
We apply Low-Rank Adaptation (LoRA) with a rank of 8 and a scaling factor ($\alpha=16$). LoRA is integrated into the following transformer layers: q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj. Models are fine-tuned for 3 epochs. The total training time for different configurations is reported in Table 9 and Table 10. The full set of hyperparameters used during training is listed in Table 11.
Hyperparameter Value
Batch Size 1
Gradient Accumulation Steps 1
16-bit Floating Point Precision True
Optimizer adamw_torch
Learning Rate $1\times 10^{-4}$
Max Gradient Norm 0.3
Warmup Ratio 0.03
LoRA Rank 8
LoRA scaling factor 16
LoRA Dropout 0.05
Weight Decay 0.01
LR Scheduler Type cosine
Table 11: Hyperparameters Used During Finetuning
D.1 Prompt for Finetuning
The prompt used for fine-tuning both Llama 3.1 8B and Mistral-Instruct-v0.2 is shown in Listing D.1.
[…]
The extracted relevant passages are inserted ahead of the "Research Paper Passages" heading. Similarly, the corresponding content for each feedback section (Summary, Strengths, Weaknesses, and Questions) is added after their respective headings in the generated feedback.
Appendix E LLM-as-a-judge and Qualitative Assessment
The prompt used for LLM-as-a-judge with Gemini-2.0-Flash is shown in Figure 4. The corresponding model-generated feedback is inserted under the appropriate subheadings within the prompt.
For Qualitative Assessment, the prompt in Figure 4 is modified only slightly as to just remove the output format. This slightly modified prompt is given as the evaluation guideline to the annotators. It is ensured that the annotators selected are eligible to review papers for the particular conference.
** Instruction ** : You are an expert on research papers. Evaluate the following feedback according to the scoring
[…]
Give rating for the categories described below on the full scale from 1(lowest rating) to 5(highest rating).
Assign scores independently per category (they may differ significantly).
**Category 1: Feedback Depth**
1 – Unqualified: Admits lack of expertise | Fundamental misunderstandings | Fails to engage with technical content
2 – Limited: Surface-level comprehension | Misses >2 key concepts | Confuses established methods
3 – Functional: Grasps main contributions | Identifies 1-2 technical flaws | Limited literature context
4 – Qualified: Precise technical critique | Contextualizes in recent work (>3 citations) |
Identifies methodological tradeoffs
5 – Authoritative: Demonstrates field leadership | Predicts follow-up research directions |
Reveals fundamental limitations in approach
**Category 2: Feedback Thoroughness**
1 – Negligent: <3 substantive comments | Skips methodology/results | Fails to address claims
2 – Patchy: Only addresses obvious aspects | 0 analysis of experiments | No ethical consideration
3 – Standard: Checks main claims | Basic methodology feedback | 1 reproducibility concern noted
4 – Rigorous: Line-by-line methodology check | Statistical validity analysis | 2+ improvement suggestions
5 – Forensic: Alternative approach proposals | Re-analysis of key results |
Compliance checklist (ethics, reproducibility, etc.)
**Category 3: Feedback Constructiveness**
1 – Destructive: Personal attacks | Unsubstantiated dismissal | 0 actionable feedback
2 – Demanding: Unrealistic requests ("redo all experiments") | Vague quality complaints
3 – Directive: General improvement areas | 1-2 specific examples | Mixed positive/negative tone
4 – Strategic: Priority-ranked suggestions | Alternative methodology paths | Conference-specific improvements
5 – Transformative: Step-by-step revision roadmap | Code/experiment snippets provided | Before-after examples
**Category 4: Feedback Helpfulness**
1 – Misleading: Factually incorrect suggestions | Contradicts paper’s goals | Harmful recommendations
2 – Abstract: "Improve writing" without examples | "Add experiments" without design
3 – Guided: Section-specific edits | 2-3 implementable suggestions | Partial error diagnosis
4 – Impactful: Directly address weaknesses | Templates/examples provided | Clear priority levels
5 – Revelatory: Exposes fundamental flaws | Provides benchmark comparisons | Supplies missing literature
**Scoring Methodology**
For EACH category:
1. Find the HIGHEST level (5-1) where ALL criteria are FULLY satisfied (Base Level)
2. Check criteria in the NEXT HIGHER LEVEL (Base Level + 1):
   - Calculate partial credit: (Satisfied Criteria in Higher Level / Total Criteria in Higher Level)
3. Final Score = Base Level + Partial Credit
4. Round to 2 decimal places (0.33 increments for 3-criteria levels, 0.5 for 2-criteria levels)
[…]
Respond ONLY with this JSON format:
  "feedback_scores": {{
    "Category 1": "",
    "Category 2": "",
    "Category 3": "",
    "Category 4": ""
Figure 4: Prompt used during evaluating the quality of generated feedbacks with LLM as a Judge
Appendix F Human Evaluation UI
Our annotation UI is developed using Streamlit  [^4] https://streamlit.io/.
F.1 Qualitative Assessment
Examples of the user interface used for the Qualitative Assessment are shown in Figure 5 and  6.
Figure 5: This interface component corresponds to the Qualitative Assessment task described in Section 6.3.1. It marks the beginning of the feedback presentation. Due to space constraints, only the summary and a subset of the strengths from one feedback are shown here.
Figure 6: This interface component corresponds to the task described in Section 6.3.1. In this part, annotators provide a final rating across the four categories and submit their annotation.
F.2 Factual Validity Assessment
Dataset Feedback Positive Rating (%)
ICLR 2024 ICT_10 76.97
Llama $(5,3)$ 65.25
Llama $(5,5)$ 64.58
Other Llama $(5,3)$ 61.86
Llama $(5,5)$ 64.31
Table 12: Overall Positive Rating Percentage (sum of “Completely Agree” and “Mostly Agree”) for test dataset of ICLR 2024 as well as NeurIPS and COLM 2025 (Other).
Examples of the user interface used for Factual Validity Assessment can be seen in Figures 7a and  7b.
(a) The original paper is displayed in a PDF viewer to provide factual context.
(b) Each point of the review is rated for accuracy using a dropdown menu.
Figure 7: The user interface for the Factual Validity Assessment task described in Section 6.3.2. On the left (a), the full research paper is presented to the annotator. On the right (b), the annotator rates each point of the feedback based on its consistency with the source document.
Appendix G Evaluation Guidelines for Factual Validity Assessment
The evaluation guidelines for the factual validity assessment are shown in Figure 8. It is ensured that the annotators selected are eligible to review papers for the particular conference.
**Instructions**:
To assess a structured feedback and check how accurately it reflects the paper’s content.
**Evaluation Process**:
1. Read the entire paper first.
   - If you need to download the paper for some reason, do ensure that you only download the paper using the provided UI.
2. Read the structured feedback and identify every DISCRETE CLAIM.
   - For the “Summary” section of the review, treat the entire content of the Summary
     as ONE discrete claim.
   - For the “Strengths”, “Weaknesses”, and “Questions” sections of the review,
     treat EACH bulleted point as a separate discrete claim.
3. For each claim, choose the best category from the dropdown.
**Claim Categorization**:
You must rate the Validity Agreement of every claim on a scale of 1–4,
where 1 indicates “Completely Disagree” and 4 indicates “Completely Agree”.
**Completely Disagree (Invalid Claim) [1]**:
- The claim is factually wrong, contradicts the paper, or is hallucinated.
**Mostly Disagree (Largely Invalid Claim) [2]**:
- The claim is mostly incorrect, though it may contain minor truth.
**Mostly Agree (Largely Valid Claim) [3]**:
- The claim is mostly correct, with minor gaps or oversights.
**Completely Agree (Fully Valid Claim) [4]**:
- The claim is entirely accurate.
Figure 8: Guidelines used for human evaluation of factual validity of the structured feedback
Appendix H Example of a Generated Structured Feedback
We present a structured feedback generated by our system in Figure 5
On the Role of Edge Dependency in Graph Generative Models
Summary
The paper introduces a novel framework for evaluating generative models of graphs,
focusing on the concept of ‘‘overlap’’ to ensure both accuracy and edge diversity.
It categorizes graph generative models into three levels of complexity: edge
independent, node independent, and fully dependent models. The paper derives
theoretical bounds on triangle counts and other short-length cycles based on overlap
and provides new generative models for each level. However, the paper’s novelty and
practical contributions are questioned due to the use of existing methods and the
lack of significant empirical validation. The experimental section is seen as
limited and does not convincingly demonstrate the superiority of the proposed
methods over existing ones.
Strengths
- The paper introduces a novel evaluation framework that considers the importance of
  model-generated graph overlap to ensure both accuracy and edge diversity, which is
  a significant contribution to the field of graph generative models.
- The paper presents a hierarchy of graph generative models categorized into three
  levels of complexity, which encapsulates a wide range of prevalent methods and
  provides a structured framework for evaluating graph generative models.
- The theoretical bounds on the number of triangles and other short-length cycles are
  well-presented and support the findings with instances demonstrating the asymptotic
  optimality of these bounds.
- The authors introduce new generative models for each of the three hierarchical
  levels, leveraging dense subgraph discovery (Gionis & Tsourakakis, 2015), which
  is a practical application of the theoretical findings.
Weaknesses
- The novelty of the paper is limited as it primarily relies on existing methods and
  theoretical frameworks, and the experimental results do not convincingly
  demonstrate the superiority of the proposed methods over existing ones.
- The empirical evaluation is weak and does not provide sufficient evidence to
  support the claims made in the paper. The experimental results are not robust and
  the paper does not provide enough comparisons with other models to establish the
  effectiveness of the proposed methods.
- The paper has several presentation issues, including unclear definitions and
  notations, which could hinder the understanding of the content for readers not
  familiar with graph generative models.
- The practical applications and the significance of the proposed methods are not
  clearly demonstrated, and the paper lacks a detailed discussion on the limitations
  and potential future work.
Questions
- Can the authors clarify the definitions and notations used in the paper,
  particularly the ones related to the overlap and triangle density, to ensure a
  clearer understanding of the content?
- How do the authors justify the novelty of their work given the reliance on
  existing methods and theoretical frameworks?
- What specific contributions does the paper make to the field of graph generative
  models, and can these be clearly articulated?
- Why was the empirical evaluation conducted in a limited manner, and how can the
  paper be improved to provide more robust and convincing experimental results?
- Could the authors discuss the practical applications of the proposed methods and
  the significance of their work in more detail?
- Are there any plans to extend the theoretical bounds to higher-order cycles, and
  how might this affect the overall conclusions drawn in the paper?
Figure 9: Output of the generated feedback by AutoRev for the paper On the Role of Edge Dependency in Graph Generative Models [^5] https://openreview.net/forum?id=LCQ7YTzgRQ